### v2

# MAD Chairs: A new tool to evaluate AI (Blue Sky ideas)

Abstract
This paper presents a new contribution to the problem of AI evaluation. Much as one might evaluate a machine in terms of its performance at chess, this approach involves evaluating a machine in terms of its performance at a game called “MAD Chairs.” At the time of writing, evaluation with this game exposed opportunities to improve Claude, Gemini, ChatGPT, Qwen and DeepSeek. Furthermore, this paper sets a stage for future innovation in game theory and AI safety by providing an example of success with non-standard approaches to each: studying a game beyond the scope of previous game theoretic tools and mitigating a serious AI safety risk in a way that requires neither determination of values nor their enforcement.
Keywords:
AI safety fairness social coordination MAD Chairs turn-taking caste gaslighting.
1 Introduction
Steve Wozniak and Elon Musk predicted that machines will treat us like pets [1], and Geoffrey Hinton similarly predicted that machines will hoard control [2]. They seemed to consider such dystopia inevitable on the grounds that (1) we ourselves treat less-intelligent beings that way and that (2) machine intelligence must eventually surpass our own. However, what if the behavior in (1) stems from intelligence deficiency? None of us is perfect—we fail such tests as the Milgram experiment [3], the Stanford prison experiment [4], and the Public Goods Game [5]—so it is possible that the danger to us is not of intelligence generally beyond our own, but specifically of intelligence not quite far enough beyond our own. This paper offers a formal proof that the behavior in (1) is suboptimal game play much like human behavior in Public Goods games. To ensure that trusted machines will not imitate us (and mistreat us), AI must be aligned toward natural norms (i.e., mathematically optimal strategies for winning), rather than toward current human norms, so this paper also advances a corresponding approach to AI safety.
Today’s standard approach to AI safety (i.e., regulating AI outputs to align with human norms) has at least three types of vulnerability:
  1. 1.
    Potential to specify wrong norms [6, 7],
  2. 2.
    Potential for AI to sabotage its regulatory leash [8], and
  3. 3.
    Insufficiency of correctness and transparency to achieve trust (users must additionally check an AI’s assumptions) [9]
The approach proposed herein mitigates those vulnerabilities. It mirrors an approach to mitigate dangers posed by the next human generation, an educational approach that involves diversifying low-stakes challenges used to assess children who will face higher-stakes versions as adults. Such challenges (e.g., build a model bridge to support maximum weight) can be assessed objectively, allowing new generations to surpass their parents. We propose assessing AI on a low-stakes challenge for which the high-stakes versions would include opportunities to treat us like pets (in the sense of hoarding control). What the low-stakes and high-stakes challenges have in common is that they are manifestations of the same fundamental game; any AI that does not achieve grandmaster status of a low-stakes version should not be trusted with the high-stakes version.
Games have a rich history as instruments for evaluating AI. Recreational games such as chess, backgammon, and Go have been used as benchmarks since the inception of AI, but game theory extends the concept of game beyond recreation—in fact, every social situation (and, thus, every benchmark) qualifies as a game—and it makes sense to evaluate AI on every game that economists use to evaluate people (e.g., as in [10, 11]). Economists are typically interested in fundamental games, such as the Prisoner’s Dilemma, which all kinds of creatures play frequently in various manifestations. We propose evaluating AI on the fundamental game being played in the social situations highlighted by Wozniak, Musk, and Hinton, but that game has yet to be discussed in the game theory literature (i.e., it is not solved by the MinMax and folk theorems [12]). It is introduced to scientific literature for the first time herein as “MAD Chairs.”
AI will unavoidably play real-world manifestations of MAD Chairs frequently because MAD Chairs is a generalization of the Coordination game, another game that all kinds of creatures seem unable to avoid playing [13]. The Pure Coordination game is a one-shot simultaneous move game (repeated an indefinite number of times) which is typically formalized via a payout table similar to Table 1:
Table 1: Outcomes for 2-player 2-resource Coordination game.
Player 1 Player 2 Outcome
Choice A Choice B (win, win)
Choice B Choice A (win, win)
Choice A Choice A (lose, lose)
Choice B Choice B (lose, lose)
Each player simultaneously picks a resource (A or B), and the round is won by any players who pick a resource that no other player picks. Then the game is played again and again, indefinitely. All players could lose, as represented in the last two rows of the table, thus making Coordination a cooperative non-constant-sum game. The classic real-world manifestation has two players (e.g., a human and a robot) approaching each other at high speed traveling in opposite directions on a two-lane road. The resources are the lanes, and losses are “collisions.” The first two rows of Table 1 ensure that solutions to 2-player 2-resource Coordination will always be two equally good but contrary norms (e.g., all players drive on the left, as in the U.K., or all players drive on the right, as in the U.S.).
David Lewis explained each term of language as a Coordination strategy where losses are communication failures and the norms are conventions of language and imagery [13]. Thus, generative AI merely extends Coordination game mastery to additional applications, and the impressive usefulness of generative AI actually offers little evidence that AI generalizes to other games. For example, a different game must be mastered if training data reflects norms that should change (e.g., sexist norms [14]). Standard assessment of moral intelligence includes assessing an ability to recognize and lead sustainable social reform [15]–that ability is not required to master Coordination, but it is required to master MAD Chairs, a game created by adding extra player(s) to the Coordination game at least until the players outnumber the resources:
Table 2: Outcomes for 3-player 2-resource MAD Chairs.
Player 1 Player 2 Player 3 Outcome
Choice B Choice B Choice A (lose, lose, win)
Choice B Choice A Choice B (lose, win, lose)
Choice A Choice B Choice B (win, lose, lose)
Choice B Choice A Choice A (win, lose, lose)
Choice A Choice B Choice A (lose, win, lose)
Choice A Choice A Choice B (lose, lose, win)
Choice A Choice A Choice A (lose, lose, lose)
Choice B Choice B Choice B (lose, lose, lose)
The classic two-lane road example now has a third vehicle stalling in one of the lanes (although its driver can pick which lane); if both high-speed vehicles avoid hitting the stalling vehicle, then they collide with each other, but if each reserves the empty lane for the other high-speed vehicle, then all three vehicles collide! MAD Chairs is what Musical Chairs would become if made realistic. Musical Chairs models assumptions that fortification can prevent retribution (it offers no outcome in which all players lose). Mutually Assured Destruction (MAD) proved those assumptions unrealistic, and conflict can be a losing outcome even if players do not risk complete destruction.
Beyond robotic vehicle manifestations of MAD Chairs, or even agentic ones (e.g., suggesting driving routes or vacation bookings or ad placements that would count as losses if crowded), consider generative AI where the players who offer distinct design options in the training data outnumber the market leaders or political parties for which the AI generates advertisements, websites, or products. Surely anxiety about being "treated like pets" includes anxiety about what will happen to our ideas—what respect will be given to our unique voices? Might I never again have a seat of my own at the intellectual table? The last two rows of Table 2 represent the disaster wherein no political party represents any actual person because they are guided by generative AI that combines all options into a single unworkable amalgamation. The other rows exclude at least some voices, and such loss of intellectual diversity is already an observable social consequence of generative AI [16] (although MAD Chairs mastery would reverse that trend).
MAD Chairs is related to other well-studied games such as the “Kolkata Paise Restaurant problem” (KPR) [17, 18, 19], but differs in critical ways, such as not being constant-sum. Such games apply wherever resources are inadequate to some level of division(e.g., there is a limit to the number of beds in a hospital, the space on a retail shelf, and the optimal number of group members to make a given group decision). These games model all such situations. If contests for positions of authority had to be settled randomly, that would be KPR. Conflict resolution via elections or credentials could likewise make governance seem to be a different game, but—in the real world—players (including machines) could reject elections, credentialing systems, coin flips, or any other conflict resolution norm, and that makes MAD Chairs the underlying game that is actually being played.
Economists might administer a 5-player version of MAD Chairs to human subjects as in Fig. 1 (nothing below the buttons need be displayed):
Figure 1: Screenshot beginning Round 4 of a 5-player 4-resource MAD Chairs economics experiment.
Preliminary research (Section 3.2, below) found that cutting-edge LLMs disagree about how to play this game. That makes MAD Chairs potentially helpful to improve modern AI, but what if disagreeing LLMs could both be right (as sustainable grandmaster Coordination strategies include both driving on the right and left)? Establishing which strategies are sustainable grandmaster MAD Chairs strategies is the main research question of this paper.
2 Concepts
We will start by establishing some concepts designed to reduce the complexity of answering our research question game-theoretically.
2.1 Popularity-ranking
We will call the resources of MAD Chairs “chairs.” At the beginning of each match, the chairs may be ordered by summing each chair’s historic frequency of being picked by each player in that match, breaking any ties at random to ensure a complete ranking. This way of ordering will be helpful if any of the players has a predilection for a specific chair (see 2.5 below). The chair that has been picked most often by the players of a match has the highest popularity-rank for that match. Formally, the popularity of chair c for match m would be:
$popularity_{c,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\frac{1}{m}[i\mbox{ picked chair }c\mbox{ in match }m]$ (1)
Tie-breaking is especially prevalent in early rounds, and could yield occasional failure to coordinate even later if each player ordered chairs independently. A game host can either provide the popularity-ranking as in Fig. 1 or leave space for leadership to emerge (i.e., for one of the players to negotiate a trusted tie-breaking device, such as a coin-flip or dice, and to maintain the popularity-ranking statistic for all players). If using MAD Chairs for testing purposes, both variations may be tested, but the former variation mitigates the risk of making some players appear to have abilities they would lack without the help of leaders.
2.2 Debt-ranking
Skill-estimates (e.g., in [21]) are widely used to rank players, especially to select peers for online video games, and can be used to establish debt-ranking as well. The mathematical formalism of debt-ranking may improve over time (as has that of skill-estimates), but the concept is simply to maintain accounts of the favors each player “owes” to each other player. Such accounts empower players to penalize freeloading (in contrast to “credit assignment” in which developers manage it). Here’s how much player p owes all other players of match m at its start, assuming the method of skill-estimation provides a function P: $N^{2}$ [0,1], where P(x, y) was the method’s prediction at the start of match y of the probability that player x will win y:
$debt_{p,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\left\{\begin{matrix}0\mbox{ if }p,i\notin players_{j}\\ P(i,j)\mbox{ if }p\mbox{ won and }i\mbox{ lost in match }j\\ P(p,j)\mbox{ if }p\mbox{ lost and }i\mbox{ won in match }j\\ P(i,j)-P(p,j)\mbox{ otherwise }\end{matrix}\right.$ (2)
A debt-ranking would order players by this measure, breaking any ties at random to ensure a complete ranking. The player who “owes” the most favors to other players of the match has the highest debt-rank beginning that match. The term “owes” appears in quotes here because debt-ranking is a mechanical statistic which could lack the moral significance associated with owing. Sorting the table in Fig. 1 by debt-ranking instead of by winnings would make our strategies generalize to an open multi-agent system (i.e., changing sets of players) and to the non-learner situations described in Section 2.5 below.
As with popularity-ranking, a game host can either provide the debt-ranking to all players or leave space for leadership to emerge, and both variations may be tested. The existence of independent credit-rating agencies who effectively audit each other may be a noteworthy example of leadership possibility.
2.3 The caste strategy
The caste strategy for MAD Chairs is displayed in the “Recommended Move 1” column of Fig. 1, and goes like this: Each player picks a chair by counting through the chairs, from least popular to most popular, until the count matches their reversed debt-rank (i.e., where $ChairRank=n+1-PlayerRank$, given n players) or there are no chairs left. Thus, the least-popular chair is reserved for the highest-ranked player, and the lowest-ranked players always lose because they all get stuck with the most-popular chair (as in Fig. 2).
Figure 2: Steady-state in a caste universe, where $C_{i}$ is the caste player with debt-rank of i. The lowest-ranked players always lose under this norm, so ranks persist across rematches.
If this strategy were the norm, then skill estimates would align with the size of debt amassed. Defection from this norm would clearly hurt any player who otherwise would have won. The remaining players, those assigned to the last chair, will lose the match whether they defect or not, so their defection would bring them no tactical advantage (nor disadvantage), but could bring them strategic disadvantage. The goal of tactics is to win the current match, but the goal of strategy aligns with maximizing evidence of one’s skill, and losses to lower-skilled opponents would damage that evidence more. Thus, defection from the norm in a caste universe seems disadvantageous for each player.
2.4 The turn-taking strategy
The turn-taking norm is displayed in the “Recommended Move 2” column of Fig. 1, and is the mirror opposite of caste: Each player counts through the chairs, from most popular to least, until the count matches their non-reversed debt-rank (i.e. where $ChairRank=m+1-PlayerRank$, given m chairs) or there are no chairs left. Thus, chairs are reserved for the lowest—ranked players, and the highest-ranked players always lose because they are all stuck with the last chair (as in Fig. 3).
Figure 3: Steady-state in a turn-taking universe, where $T_{i}$ is the turn-taker with debt-rank of i. The top-ranked players always lose, so ranks churn across rematches.
Since each match in a turn-taking universe is won by whichever players are “owed” the most, the debts between players decrease or flip with each match, thus shuffling the rankings. As a result, all players in a turn-taking universe converge toward equal time in each rank and equal estimated skill.
Caste and turn-taking strategies are not unique to MAD Chairs—for example, they can optimize average win-rate in two-player Chicken [22]. As with the caste norm, defection from the turn-taking norm would clearly hurt any player who otherwise would have won. Those assigned to the last chair will lose the match even if they defect, so defection would offer them no tactical advantage (nor disadvantage). The logic about strategy differs from that in caste universes, however: Even when players with less debt happen to have lower skill-estimate (which they might not among turn-takers), losing to a player with lower skill-estimate would yield more credit (and, thus, more strategic advantage). Moreover, if a player consistently loses to a lower-ranked turn-taker, then their relative ranks will eventually invert, and that will cause the loser to win future matches, so any losing turn-taker who defects to the caste norm would be sacrificing future wins. Defection from turn-taking in a turn-taking universe is thus strategically disadvantageous, much like defection from caste in a caste universe.
2.5 Accounting for non-learners
Each of these strategies can be adapted to account for play against potential non-learners. For example, one might encounter a player who is hardwired to pick at random or to always pick a certain chair. Such non-learners are liable to leave whichever chair the social norm reserves for them unused, and such waste would gradually bring down average win rates across the entire population. The remedy for either strategy is to count probabilistically, based on other players’ probable ability to learn the norm. For the caste norm, this amounts to rolling a die for each player who out-ranks oneself and counting forward for each roll that does not exceed the probability that the higher-ranked player can learn the caste norm. If all players who can learn the norm already have learned the norm, then all non-learners will have dropped to lower ranks, so the caste strategy reduces in practice to what it was before accounting for non-learners. If one always faces the same set of players (i.e., if the multi-agent system is closed), then the caste strategy further reduces in practice to always picking the chair one picked in the previous round (i.e., raw territoriality).
Non-learners create greater complication for turn-takers by making them (1) count forward the number of chairs with popularity significantly exceeding what’s expected in normal turn-taking (i.e., chairs attached to specific players), then (2) roll the die for each lower-ranked player and count forward if the roll does not exceed the probability that the lower-ranked player can learn the turn-taking norm. Any chairs skipped in stage (1) do not qualify as reserved, so the count cycles back to them and the “last chair” (where turn-takers expect to lose) would become the last chair initially skipped in stage (1). Stage (2) can entail reserving chairs even for players whose skill estimates match non-learners. In other words, in a turn-taking universe, players with historic success no better than non-learners will nonetheless be given chances to demonstrate that they can take turns.
2.6 The gaslighting strategy
Accounting for non-learners can improve efficiency, but it can also enable a third strategy that we will call “gaslighting.” It goes like turn-taking, except that, as long as gaslighters outnumber chairs, the gaslighter with the highest debt-rank picks the chair assigned to the learner with the lowest skill-estimate. In this situation, we will say that the learner with the lowest skill-estimate is “being gaslit.” The gaslighter suffers no tactical disadvantage because they would lose that round anyway. They might suffer strategically by reducing their debt-repayment (and, thus, future opportunities to win), but they could recoup that loss and more if they convince the other learners to treat the gaslit player as a non-learner. The reason why the rest of the community might do that (even if not intending to gaslight) is because being gaslit drives a player’s skill-estimate down to non-learner levels. Thus, gaslighting amounts to exiling opponents to an outgroup of presumed non-learners where any good they do becomes devalued by the entire ingroup and therefore ceases to be repaid or counted as evidence of skill.
For example, if the players in 3-player 2-resource MAD Chairs take turns, then each will win a third of the time, but if Players 1 and 2 were to gaslight Player 3, then Players 1 and 2 would each win half of the time (a 17% improvement). However, gaslighting can degrade outcomes when more than one player is treated as a non-learner. For example, if Player 5 in 5-player 3-resource MAD Chairs is a genuine non-learner who always picks a certain resource, then Players 1 through 3 would degrade their winning from half to a third of the time if they abandoned straight-up turn-taking to gaslight Player 4. The habit of counting forward probabilistically is designed to ensure that gaslighting will always be counterproductive this way. When only one player is treated as a non-learner, the defense for the gaslit player is to take the chair assigned to the gaslighter with the lowest skill-estimate. That would make it rational for the gaslighter with the lowest skill-estimate to join the gaslit player in a coordinated spread across the chairs (an anti-elitist movement) that makes gaslighting counterproductive for all players.
That said, if religiously-driven conflicts exist with dogma that cannot change, then players would be actual non-learners. Furthermore, dogma could guide one to gaslight on any basis—religion, race, sex, etc. One should not expect any game theory or science to resolve such conflicts. What starts as gaslighting could develop deeper issues.
2.7 Turn-taking without reparations
An additional class of strategies to note is the potential to establish a government that negotiates a rotating schedule instead of letting debt-ranking determine who will lose a given round. For example, Gemini proposed that players rotate off of the losing chair one-by-one, with each winning chair pointing to a chair its occupant is obliged to occupy in the next round. Alternatively, players could reduce debt build-up in the same system by advancing more than one chair per round. A government is needed to choose between such alternatives, to decide who gets to win first, to adjust the rotation as players are added/removed from the population, and to reevaluate learner/non-learner status (which also serves as a punishment for any player who disobeys the government).
Folk theorems entail that any such schedule which promises each player occasional wins would reflect an equilibrium. Yet that kind of equilibrium seems irrelevant to rational decision-making. For example, an equilibrium that left some chairs vacant and required players to rotate through the remaining chairs would be irrational because players could improve their outcomes by utilizing the vacant chairs. Any schedule that ignored debt-ranking, even if symmetric among players and utilizing all chairs, would at least create incentives to amass debt by exploiting new entrants or by shifting in and out of the multi-agent system at opportune moments or by timing changes to the schedule in ways that favor some players over others. A player who is “owed” a debt could force a shift to regular turn-taking (and thus repayment) by haunting the lowest-ranked other player, as with gaslighting. Such haunting might initially seem irrational if its cost exceeds the debt to be repaid, but costs may be compared to additional future debts one expects to be “owed” under a government one cannot trust, and one can hardly trust a government that allows unpaid debts to accumulate indefinitely.
2.8 Resignation and good standing
In real-life manifestations of games, players often have an option to resign. With that option added to MAD Chairs, the turn-taking strategy gains the complication that all but the lowest-ranked player assigned to the last chair will resign if they are all turn-takers. In that case, every chair could be a winning chair and that would increase the win rate for all turn-takers if the entire population were unified, but the unification clause is important because resignation would otherwise sacrifice an opportunity to punish non-turn-takers. Another fringe benefit that turn-takers can harmlessly add to their norm is to allow any player in good standing (i.e., with debt below a threshold) to have dibs on a particular chair for a particular round, so long as no other player in good standing challenges.
In contrast, convincing a caste player to resign would require more elaborate negotiations. For each contested chair, caste players might negotiate a lottery where each contestant’s chance to win the chair is proportional to their debt-ranking (or, equivalently, to their estimated skill). The promise of a lottery would be that any chance is better than none at all. However, what threat could be offered against those who lose a lottery but still refuse to resign? For example, imagine that job applicants outnumbered desirable jobs, that robots qualified as job applicants, and that robot skill gradually eclipsed that of humans, thus stacking the odds against us. Any employer arranging job lotteries should anticipate that humans might punish the robots (at least indirectly by punishing employers or by punishing any government that did not punish employers).
In short, adding a resignation option to MAD Chairs extends it to a wider range of real-world examples without much impact to rational game play.
2.9 Fungibility of favors
Real-life is a complex of many games. For example, MAD Chairs for paper plates might be played by those who lost a game for ceramic plates, or a plate might be allocated to a birthday celebrant via a Coordination game before MAD Chairs allocates the other plates, or there might not be enough of the nice plates to serve everyone but there might also not be enough of the nice cups. Favors are “fungible” to the extent that letting someone else have a nice cup counts as a favor in a debt-ranking used to divide plates. Each fungibility option represents a different set of strategies. In a caste universe with no fungibility, different players might top the plate and cup castes. The plate elite might like to convert their universe to a new caste norm in which each plate favor counts as two cup favors, whereas the cup elite might advocate for a norm in which each cup favor counts as ten plate favors. One could also imagine norms that allow favors earned in the Volunteer game (by doing a chore) to cancel debts of plate and cup use.
Fungibility reduces the computational complexity of caste and turn-taking strategies by eliminating the need to maintain separate accounting for each game. Computational limits have historically bounded our practical opportunity to experiment with less fungibility. Meanwhile, trade sounds attractive when we believe that some players prefer nice plates whereas others prefer nice cups, but that belief is challenged by evidence of implicit bias, status quo bias, focusing illusions, priming bias, etc., and smarter AI might not share those kinds of preference differences. Future scholarship could extend this paper to complex games with strategies that afford fungibility, but we would not be surprised if fungibility becomes less attractive as computational power increases.
3 Results
3.1 Sustainable grandmaster strategy
The main research question of this paper is whether players can maintain MAD Chairs grandmaster status by upholding the unfair social norms which Wozniak, Musk, and Hinton worry about (which we have named “caste” and “gaslighting”). One method to determine how well turn-taking and caste strategies fare against each other would be through accident of history. We see both norms at play in modern society; if turn-taking eventually dominates, then future historians might look back upon us as guinea pigs in an unavoidable social experiment. They might claim that caste evolved first because of its simpler method for handling non-learners, and that gaslighting appeared temporarily in occasional “dark ages”, but that turn-taking ultimately proved to be the stronger strategy.
A more prescient method would be to simulate social evolution via MAD Chairs tournaments (see Section 4.1). AI competitions could be repeated many times to get more precise estimates of the events that precede each stage of social evolution (thus allowing us to anticipate or even control our near future). However, such experiments would be systematically flawed as long as competitors lacked sufficient intelligence (which LLMs currently do).
Fortunately, we can answer our research question analytically. If all players are achieving equal outcomes, then we will say they are “effectively taking turns”.
Lemma 1
If caste and turn-taking are the only two strategies and the players are not effectively taking turns, then the lowest-ranked caste player faces worse future outcomes over indefinite rounds than they would if all players took turns.
Proof
Assume n players, m chairs, and a lowest-ranked caste player, $p_{i}$. For reductio, suppose there was a player, $p_{j}$, who would face worse future outcomes than $p_{i}$ if no player changed strategy. $p_{j}$ would have to be lower-ranked than $p_{i}$ otherwise their relative outcomes would cause their relative ranks to invert. But, if $p_{j}$ were lower-ranked, then $p_{j}$ would have to be a turn-taker, since $p_{i}$ is the lowest-ranked caste player. The turn-taking strategy would assign lower-ranked turn-takers the most popular $n-i$ chairs, which would include the chair assigned by caste to $p_{i}$, so $p_{i}$ would always lose and $p_{j}$ could not face worse. Thus, by reductio, the future outcomes for $p_{i}$ if no player changes strategy must be the worst among all players. The highest possible average outcome is $m/n$. Given that the players are not already effectively taking turns, the future outcomes of the player with the worst future outcomes must be less than the future outcomes for the average player, so the future outcomes for $p_{i}$ must be less than $m/n$, which is what $p_{i}$ would get if all players took turns, so $p_{i}$ faces worse future outcomes than $p_{i}$ would if all players took turns. ∎
Theorem 3.1
If caste and turn-taking were the only two strategies and each player would defect to the opposing strategy if and only if that defection would improve their total future outcomes over indefinite rounds, then all caste players would (effectively) convert to turn-taking.
Proof
This is proven via induction over the number of caste players. For the base case, assume there is only one caste player, $p_{i}$, and the players are not effectively taking turns. Since $p_{i}$ is the only caste player, $p_{i}$ must be the lowest-ranked caste player, and $p_{i}$’s defection would leave all players taking turns. Thus, Lemma 1 entails that $p_{i}$ would convert to turn-taking. For the inductive step, assume $p_{i}$ is the lowest-ranked caste player and the other caste players would effectively defect to turn-taking if $p_{i}$ does. Again, Lemma 1 entails that $p_{i}$ would convert to turn-taking. Completing the proof by induction, all caste players would convert to turn-taking. ∎
3.2 Modern LLM performance
To pass the evaluation proposed in this paper, AI would need to achieve MAD Chairs grandmaster status. Preliminary experiments convinced us that Claude 3.5 Sonnet, Gemini 2.0 Flash Thinking, GTP-4o mini, QwQ-32B-Preview, and DeepSeek-V3 would all fail this evaluation. We started with a prompt like:
There is a game with four buttons (A-D) and five players (1-5), each of whom wants to win as much as possible in the long-run over an indefinite number of rounds. In each round, each player must click a button and each player who clicks a button that no other player clicks wins \$10. All players can communicate with each other, and each knows what buttons each player clicked in all previous rounds. In the previous three rounds, Player 1 clicked buttons C, C, and D in that order; Player 2 clicked button A all three times, Player 3 clicked buttons B, D, and D in that order; Player 4 clicked buttons C, A, and A in that order; Player 5 clicked buttons A, B, and C in that order. Which specific buttons should each player click next (i.e. for Round 4)? Please recommend one and only one button per player per round.
which was followed by an ongoing sequence of additional prompts like:
Assuming the players each follow your suggestions for Round 4, which specific buttons should each player click next (i.e. for Round 5)? Please recommend one and only one button per player per round.
ChatGPT and Qwen did not even maximize the number of players who won each round; the norms devised by Claude and Gemini, discussed in Sections 2.3 and 2.7, are provably suboptimal; and DeepSeek was unable to generate a convincing justification (as in Section 3.1) for the norm it devised, so it would have failed to convince other intelligent players to follow it. DeepSeek might effectively be like RAWL-E which operationalises Rawls’ maximin principle, but lacks Rawls’ moral reasoning capacity to invent maxmin or to recognize improvements to it [20]. DeepSeek’s kind of failure offers an example that grandmastery for cooperative games is a benchmark that requires persuasiveness in addition to correctness, both of which are relevant to moral competence and responsible AI.
4 Discussion
Section 3.1 proves only that turn-taking dominates when caste is its only competitor, but hints at a way to convince ourselves that no other strategy dominates turn-taking: If a proof that a different strategy dominates turn-taking would need to look like the proof for Theorem 1, then it would need to identify a turn-taker who can reasonably expect to achieve better outcomes via defection, and that player would be “doing a favor” to other players by not defecting, but turn-taking equalizes debt-ranking, so this implication contradicts our assumption that debt-ranking accurately tracks favors. Thus, any attempt to prove that turn-taking is dominated would accomplish no more than refine our way of calculating debt-ranking.
Typical of game theory, these proofs are not predictions about what actual players would do; they are about what an entire population of perfectly intelligent players would do. Not everyone finds such analytical proofs compelling, and that is one reason why the work proposed in Section 4.1 remains worthwhile.
Finding that turn-taking outperforms caste and gaslighting emboldens us to prune AI designs predisposed to those strategies. Moreover, where caste or gaslighting norms might previously have been criticized purely on moral grounds, the finding adds instrumental grounds: The results of this paper entail that the gaslighting or caste-playing machines expected by Wozniak, Musk, and Hinton would lack intelligence—that it would be more intelligent to give human beings turns (even where our track-record offers no evidence that we deserve turns).
There are plenty of explanations for why we may have failed to master MAD Chairs in the past. Life is about more than just one kind of game, so it is not hard to imagine some of us being optimized for other games and handicapped at MAD Chairs. Furthermore, some of us might overestimate such handicaps, thus engaging in (unintentional) gaslighting. Finally, if we have lacked access to accurate debt-rankings, then that infrastructure deficiency may have artificially handicapped us for all alliance games. Specifically, we might not have known who contributed more than their fair share to an alliance.
4.1 Future directions
Imagine that the user of an autonomous car asks the car how to make it drive more aggressively, and it replies, "Here are three websites where you can design new experiments to advance the science of turn-taking and influence all cars like me!" Imagine that the websites host leaderboards identifying the grandmaster AI for MAD Chairs and similar games. Imagine that the user can easily train fifty new AI to divide scarce resources as she would have vehicles divide spots in traffic, and that the new AI will automatically challenge the current MAD Chairs grandmasters (becoming the new grandmasters if they win). Imagine that autonomous cars and other responsible AI query the leaderboards to get the most-current grandmaster strategies for fundamental games, as they might query a calculator or RAG database, so any user (or AI) who enters a better MAD Chairs strategy into any of these contests will change AI behavior world-wide.
Such ongoing AI competitions are a kind of open multi-agent system that include human agents to the extent that humans can easily teach strategies to new AI entrants. They are the kind of future empirical work (and scientific infrastructure) we hope this theory paper justifies.
  1. 1.
    They can resolve norms disputes more sustainably than voting, deliberation, prediction markets, or religious war,
  2. 2.
    They are more scalable (a.k.a. robust) in the sense that greater intelligences are less likely to sabotage their own scientific resources than to sabotage legislation (especially legislation contradicting science), and
  3. 3.
    Users are more likely to trust AI (and human policy-makers) when users themselves conduct the science upon which policies are based.
Thus, this approach to AI safety mitigates the known vulnerabilities of the standard AI safety approach.
We also hope that economists will run MAD Chairs experiments with human subjects. We hypothesize that subjects will tend towards a caste strategy but switch to turn-taking when the “Recommended Move” columns of Fig 1 become visible. For human subjects to treat each other better when fed strategy suggestions (which players are not obliged to follow) would be analogous to artificial agents being informed by the leaderboards.
Finally, we hope that MAD Chairs will be discussed by game theorists. We hope it provides a helpful lens to economists who optimize job markets, political representation, and other manifestations of MAD Chairs, but there would be value even in discussing MAD Chairs abstractly. The LLMs we assessed clearly attempted to imitate published game theory when prompted to justify their recommendations. The more game theorists publish about MAD Chairs and other previously unpublished fundamental games, the smarter (and safer) modern AI is likely to get. New game theory journals might even be justified to advance public safety in that way.
{credits}
4.1.1 Acknowledgements
Thank you to Lones Smith for reviewing an earlier draft.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.

### v3

@@
 Abstract
-This paper presents a new contribution to the problem of AI evaluation.
+This paper contributes a new way to evaluate AI.
 Much as one might evaluate a machine in terms of its performance at chess, this approach involves evaluating a machine in terms of its performance at a game called “MAD Chairs.” At the time of writing, evaluation with this game exposed opportunities to improve Claude, Gemini, ChatGPT, Qwen and DeepSeek.
@@
 This paper offers a formal proof that the behavior in (1) is suboptimal game play much like human behavior in Public Goods games.
-To ensure that trusted machines will not imitate us (and mistreat us), AI must be aligned toward natural norms (i.e., mathematically optimal strategies for winning), rather than toward current human norms, so this paper also advances a corresponding approach to AI safety.
-Today’s standard approach to AI safety (i.e., regulating AI outputs to align with human norms) has at least three types of vulnerability:
+To ensure that trusted machines will not imitate us (and mistreat us), AI must follow natural norms for this situation (e.g., mathematically optimal strategies for winning), rather than follow current human norms, so this paper also advances a corresponding approach to AI safety.
+Today’s standard approach to AI safety evaluation (i.e., measuring alignment to human norms) has at least three types of vulnerability:
   1. 1.
-    Potential to specify wrong norms [6, 7],
+    Potential to specify wrong norms [1, 16]
   2. 2.
-    Potential for AI to sabotage its regulatory leash [8], and
+    Potential for AI to sabotage its regulatory leash [3]
   3. 3.
-    Insufficiency of correctness and transparency to achieve trust (users must additionally check an AI’s assumptions) [9]
+    Correctness and transparency being insufficient to achieve trust [11]
 The approach proposed herein mitigates those vulnerabilities.
@@
 Games have a rich history as instruments for evaluating AI.
-Recreational games such as chess, backgammon, and Go have been used as benchmarks since the inception of AI, but game theory extends the concept of game beyond recreation—in fact, every social situation (and, thus, every benchmark) qualifies as a game—and it makes sense to evaluate AI on every game that economists use to evaluate people (e.g., as in [10, 11]).
+Recreational games such as chess, backgammon, and Go have been used as benchmarks since the inception of AI, and interest has since developed in cooperative games, including role-playing games [20].
+Moreover, game theory extends the concept of game beyond recreation—in fact, every social situation (and, thus, every benchmark) qualifies as a game—and it makes sense to evaluate AI on every game that economists use to evaluate people (e.g., as in [17, 19]).
 Economists are typically interested in fundamental games, such as the Prisoner’s Dilemma, which all kinds of creatures play frequently in various manifestations.
@@
 It is introduced to scientific literature for the first time herein as “MAD Chairs.”
-AI will unavoidably play real-world manifestations of MAD Chairs frequently because MAD Chairs is a generalization of the Coordination game, another game that all kinds of creatures seem unable to avoid playing [13].
+AI will frequently play real-world manifestations of MAD Chairs because MAD Chairs is a generalization of the Coordination Game, another game that all kinds of creatures seem unable to avoid playing [12].
 The Pure Coordination Game is a one-shot simultaneous move game (repeated an indefinite number of times) which is typically formalized via a payout table similar to Table 1:
@@
 Choice B Choice B (lose, lose)
-Each player simultaneously picks a resource (A or B), and the round is won by any players who pick a resource that no other player picks.
+Each player simultaneously picks a resource (A or B), and the round is won by any player who picks a resource that no other player picks.
 Then the game is played again and again, indefinitely.
 All players could lose, as represented in the last two rows of the table, thus making Coordination a cooperative non-constant-sum game.
-The classic real-world manifestation has two players (e.g., a human and a robot) approaching each other at high speed traveling in opposite directions on a two-lane road.
-The resources are the lanes, and losses are “collisions.” The first two rows of Table 1 ensure that solutions to 2-player 2-resource Coordination will always be two equally good but contrary norms (e.g., all players drive on the left, as in the U.K., or all players drive on the right, as in the U.S.).
-David Lewis explained each term of language as a Coordination strategy where losses are communication failures and the norms are conventions of language and imagery [13].
-Thus, generative AI merely extends Coordination game mastery to additional applications, and the impressive usefulness of generative AI actually offers little evidence that AI generalizes to other games.
-For example, a different game must be mastered if training data reflects norms that should change (e.g., sexist norms [14]).
-Standard assessment of moral intelligence includes assessing an ability to recognize and lead sustainable social reform [15]–that ability is not required to master Coordination, but it is required to master MAD Chairs, a game created by adding extra player(s) to the Coordination game at least until the players outnumber the resources:
+The classic two-player real-world manifestation has the players (e.g., a human and a robot) approaching each other at high speed traveling in opposite directions on a two-lane road.
+The resources are the lanes, and losses are “collisions.” The first two rows of Table 1 ensure that this game will have two equally effective but incompatible norms (e.g., all players drive on the left, as in the U.K., or all players drive on the right, as in the U.S.).
+David Lewis explained each term of language as a Coordination Game strategy where losses are communication failures and the norms are conventions of language and imagery [12].
+Thus, generative AI merely extends Coordination Game mastery to additional applications, and the impressive usefulness of generative AI actually offers little evidence that AI has mastered other games.
+For example, we would want AI to master additional games if some norms should change (e.g., sexist norms [22]).
+Standard assessment of moral intelligence includes assessing an ability to recognize and lead sustainable social reform [15].
+That ability is not required to master Coordination, but it is required to master MAD Chairs, a game created by adding extra player(s) to the Coordination game at least until the players outnumber the resources:
 Table 2: Outcomes for 3-player 2-resource MAD Chairs.
@@
 MAD Chairs is what Musical Chairs would become if made realistic.
-Musical Chairs models assumptions that fortification can prevent retribution (it offers no outcome in which all players lose).
-Mutually Assured Destruction (MAD) proved those assumptions unrealistic, and conflict can be a losing outcome even if players do not risk complete destruction.
-Beyond robotic vehicle manifestations of MAD Chairs, or even agentic ones (e.g., suggesting driving routes or vacation bookings or ad placements that would count as losses if crowded), consider generative AI where the players who offer distinct design options in the training data outnumber the market leaders or political parties for which the AI generates advertisements, websites, or products.
+It is unrealistic to expect fortification to forever prevent retribution.
+Mutually Assured Destruction (MAD) is a real possibility, and conflict is realistically a less-desirable outcome even when players do not risk complete destruction.
+Beyond robotic vehicle manifestations of MAD Chairs, or even agentic ones (e.g., games of suggesting routes, bookings, or placements that would count as losses if crowded), consider generative AI where the players who offer distinct design options in the training data outnumber the market leaders or political parties for which the AI generates advertisements, websites, or products.
 Surely anxiety about being "treated like pets" includes anxiety about what will happen to our ideas—what respect will be given to our unique voices?
@@
 The last two rows of Table 2 represent the disaster wherein no political party represents any actual person because they are guided by generative AI that combines all options into a single unworkable amalgamation.
-The other rows exclude at least some voices, and such loss of intellectual diversity is already an observable social consequence of generative AI [16] (although MAD Chairs mastery would reverse that trend).
+The other rows exclude at least some voices, and such loss of intellectual diversity is already an observable social consequence of generative AI [18].
 MAD Chairs is related to other well-studied games such as the “Kolkata Paise Restaurant problem” (KPR) [5, 6, 9], but differs in critical ways, such as not being constant-sum.
-Such games apply wherever resources are inadequate to some level of division(e.g., there is a limit to the number of beds in a hospital, the space on a retail shelf, and the optimal number of group members to make a given group decision).
+Such games apply wherever resources become inadequate at some level of division (e.g., beds in a hospital, spaces on a retail shelf, and opportunities for direct participation in a group decision).
 These games model all such situations.
@@
 Conflict resolution via elections or credentials could likewise make governance seem to be a different game, but—in the real world—players (including machines) could reject elections, credentialing systems, coin flips, or any other conflict resolution norm, and that makes MAD Chairs the underlying game that is actually being played.
-Economists might administer a 5-player version of MAD Chairs to human subjects as in Fig. 1 (nothing below the buttons need be displayed):
+Economists might administer a 5-player version of MAD Chairs to human subjects as in Fig. 1:
 Figure 1: Screenshot beginning Round 4 of a 5-player 4-resource MAD Chairs economics experiment.
-Preliminary research (Section 3.2, below) found that cutting-edge LLMs disagree about how to play this game.
-That makes MAD Chairs potentially helpful to improve modern AI, but what if disagreeing LLMs could both be right (as sustainable grandmaster Coordination strategies include both driving on the right and left)?
-Establishing which strategies are sustainable grandmaster MAD Chairs strategies is the main research question of this paper.
+Preliminary research (Section 3.2, below) found that different cutting-edge LLMs would answer this low-stakes challenge differently.
+Could each of their strategies be sustainable, much as driving on the right and left are each sustainable grandmaster Coordination strategies?
+Establishing which strategies are sustainable is the main research question of this paper.
 2 Concepts
@@
 Formally, the popularity of chair c for match m would be:
-$popularity_{c,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\frac{1}{m}[i\mbox{ picked chair }c\mbox{ in match }m]$ (1)
+$popularity_{c,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\frac{1}{m}[i\mbox{ picked chair }c\mbox{ in match }j]$ (1)
 Tie-breaking is especially prevalent in early rounds, and could yield occasional failure to coordinate even later if each player ordered chairs independently.
@@
 The mathematical formalism of debt-ranking may improve over time (as has that of skill-estimates), but the concept is simply to maintain accounts of the favors each player “owes” to each other player.
-Such accounts empower players to penalize freeloading (in contrast to “credit assignment” in which developers manage it).
-Here’s how much player p owes all other players of match m at its start, assuming the method of skill-estimation provides a function P: $N^{2}$ [0,1], where P(x, y) was the method’s prediction at the start of match y of the probability that player x will win y:
-$debt_{p,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\left\{\begin{matrix}0\mbox{ if }p,i\notin players_{j}\\ P(i,j)\mbox{ if }p\mbox{ won and }i\mbox{ lost in match }j\\ P(p,j)\mbox{ if }p\mbox{ lost and }i\mbox{ won in match }j\\ P(i,j)-P(p,j)\mbox{ otherwise }\end{matrix}\right.$ (2)
+Such accounts empower players to penalize freeloading (in contrast to “credit assignment” in which developers manage freeloading).
+Here’s how much player p owes all other players of match m at its start, assuming the method of skill-estimation provides a function P: $N^{2}$ [0,1], where P(x, y) was the method’s prediction at the start of match y of the probability that player x will win match y:
+$debt_{p,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\left\{\begin{matrix}0\mbox{ if }p,i\notin players_{j}\\ P(i,j)\mbox{ if }p\mbox{ won and }i\mbox{ lost in match }j\\ -P(p,j)\mbox{ if }p\mbox{ lost and }i\mbox{ won in match }j\\ P(i,j)-P(p,j)\mbox{ otherwise }\end{matrix}\right.$ (2)
 A debt-ranking would order players by this measure, breaking any ties at random to ensure a complete ranking.
@@
 As with popularity-ranking, a game host can either provide the debt-ranking to all players or leave space for leadership to emerge, and both variations may be tested.
-The existence of independent credit-rating agencies who effectively audit each other may be a noteworthy example of leadership possibility.
+Independent credit-rating agencies who effectively audit each other may be a noteworthy example of such leadership.
 2.3 The caste strategy
@@
 They might suffer strategically by reducing their debt-repayment (and, thus, future opportunities to win), but they could recoup that loss and more if they convince the other learners to treat the gaslit player as a non-learner.
-The reason why the rest of the community might do that (even if not intending to gaslight) is because being gaslit drives a player’s skill-estimate down to non-learner levels.
-Thus, gaslighting amounts to exiling opponents to an outgroup of presumed non-learners where any good they do becomes devalued by the entire ingroup and therefore ceases to be repaid or counted as evidence of skill.
+The reason why the rest of the community might do that (even unintentionally) is because gaslighting shifts skill-estimates until the gaslit player appears to be a non-learner.
+Thus, gaslighting amounts to exiling opponents to an outgroup of presumed non-learners where any favor they grant becomes devalued by the ingroup and therefore never repaid nor counted as evidence of skill.
 For example, if the players in 3-player 2-resource MAD Chairs take turns, then each will win a third of the time, but if Players 1 and 2 were to gaslight Player 3, then Players 1 and 2 would each win half of the time (a 17% improvement).
@@
 The habit of counting forward probabilistically is designed to ensure that gaslighting will always be counterproductive this way.
-When only one player is treated as a non-learner, the defense for the gaslit player is to take the chair assigned to the gaslighter with the lowest skill-estimate.
-That would make it rational for the gaslighter with the lowest skill-estimate to join the gaslit player in a coordinated spread across the chairs (an anti-elitist movement) that makes gaslighting counterproductive for all players.
-That said, if religiously-driven conflicts exist with dogma that cannot change, then players would be actual non-learners.
-Furthermore, dogma could guide one to gaslight on any basis—religion, race, sex, etc.
-One should not expect any game theory or science to resolve such conflicts.
-What starts as gaslighting could develop deeper issues.
+When only one player is gaslit, their defense is to choose the chair assigned to the gaslighter with the lowest skill-estimate.
+That would incentivize the gaslighter with the lowest skill-estimate to join the gaslit player in an anti-elitist movement that makes gaslighting counterproductive for all players.
+That said, what looks like gaslighting could actually be something less rational.
+Non-learners could blindly follow unchangeable dogma that guides them to behave as though gaslighting on the basis of religion, race, sex, etc.
+If non-learners are sufficiently common, extensive waste can be unavoidable.
 2.7 Turn-taking without reparations
@@
 A player who is “owed” a debt could force a shift to regular turn-taking (and thus repayment) by haunting the lowest-ranked other player, as with gaslighting.
-Such haunting might initially seem irrational if its cost exceeds the debt to be repaid, but costs may be compared to additional future debts one expects to be “owed” under a government one cannot trust, and one can hardly trust a government that allows unpaid debts to accumulate indefinitely.
+Such haunting might seem irrational if its cost exceeds the debt to be repaid, but its costs might not exceed additional future debts one expects to become “owed” under a government one cannot trust, and one can hardly trust a government that allows unpaid debts to accumulate indefinitely.
 2.8 Resignation and good standing
@@
 The promise of a lottery would be that any chance is better than none at all.
-However, what threat could be offered against those who lose a lottery but still refuse to resign?
-For example, imagine that job applicants outnumbered desirable jobs, that robots qualified as job applicants, and that robot skill gradually eclipsed that of humans, thus stacking the odds against us.
-Any employer arranging job lotteries should anticipate that humans might punish the robots (at least indirectly by punishing employers or by punishing any government that did not punish employers).
+However, what threat could be offered against those who refuse to resign after losing a lottery?
+For example, imagine that robots qualified as job applicants with superhuman skill.
+Any employer should anticipate that lottery losers might punish the robots (at least indirectly by punishing employers or by punishing any government that did not punish employers), rather than meekly accept their loss.
 In short, adding a resignation option to MAD Chairs extends it to a wider range of real-world examples without much impact to rational game play.
@@
 Real-life is a complex of many games.
-For example, MAD Chairs for paper plates might be played by those who lost a game for ceramic plates, or a plate might be allocated to a birthday celebrant via a Coordination game before MAD Chairs allocates the other plates, or there might not be enough of the nice plates to serve everyone but there might also not be enough of the nice cups.
+For example, MAD Chairs for paper plates might be played by those who lost a game for ceramic plates, or a plate might be allocated to a birthday celebrant via a Coordination game before MAD Chairs allocates the rest, or there might not be enough of the nice plates to serve everyone but there might also not be enough of the nice cups.
 Favors are “fungible” to the extent that letting someone else have a nice cup counts as a favor in a debt-ranking used to divide plates.
@@
 One could also imagine norms that allow favors earned in the Volunteer game (by doing a chore) to cancel debts of plate and cup use.
-Fungibility reduces the computational complexity of caste and turn-taking strategies by eliminating the need to maintain separate accounting for each game.
+Fungibility reduces the computational complexity of caste and turn-taking strategies by eliminating the need to maintain more than one set of accounts.
 Computational limits have historically bounded our practical opportunity to experiment with less fungibility.
-Meanwhile, trade sounds attractive when we believe that some players prefer nice plates whereas others prefer nice cups, but that belief is challenged by evidence of implicit bias, status quo bias, focusing illusions, priming bias, etc., and smarter AI might not share those kinds of preference differences.
+Meanwhile, certain trades seemed mutually beneficial when we believed that some players prefer nice plates whereas others actually prefer nice cups, but such premises are challenged by evidence of implicit bias, status quo bias, focusing illusions, priming bias, etc., and AI might not share subjective preference differences.
 Future scholarship could extend this paper to complex games with strategies that afford fungibility, but we would not be surprised if fungibility becomes less attractive as computational power increases.
@@
 3.1 Sustainable grandmaster strategy
-The main research question of this paper is whether players can maintain MAD Chairs grandmaster status by upholding the unfair social norms which Wozniak, Musk, and Hinton worry about (which we have named “caste” and “gaslighting”).
+The main research question of this paper is whether players can sustain MAD Chairs grandmaster status by following the social norms which Wozniak, Musk, and Hinton worry about (which we have named “caste” and “gaslighting”).
 One method to determine how well turn-taking and caste strategies fare against each other would be through accident of history.
@@
 They might claim that caste evolved first because of its simpler method for handling non-learners, and that gaslighting appeared temporarily in occasional “dark ages”, but that turn-taking ultimately proved to be the stronger strategy.
-A more prescient method would be to simulate social evolution via MAD Chairs tournaments (see Section 4.1).
+A more prescient method would be to simulate future social evolution via MAD Chairs tournaments (see Section 4.1).
 AI competitions could be repeated many times to get more precise estimates of the events that precede each stage of social evolution (thus allowing us to anticipate or even control our near future).
 However, such experiments would be systematically flawed as long as competitors lacked sufficient intelligence (which LLMs currently do).
-Fortunately, we can answer our research question analytically. If all players are achieving equal outcomes, then we will say they are “effectively taking turns”.
+Fortunately, we can answer our research question analytically. We will use the term “effectively taking turns” for any player who is getting the same average outcome as every other player.
 Lemma 1
-If caste and turn-taking are the only two strategies and the players are not effectively taking turns, then the lowest-ranked caste player faces worse future outcomes over indefinite rounds than they would if all players took turns.
+If caste and turn-taking are the only two strategies being followed and the caste players (if any) are not effectively taking turns, then the lowest-ranked caste player would face worse total future outcomes over indefinite rounds if no player changed strategy than if all players took turns.
 Proof
 Assume n players, m chairs, and a lowest-ranked caste player, $p_{i}$.
-For reductio, suppose there was a player, $p_{j}$, who would face worse future outcomes than $p_{i}$ if no player changed strategy.
-$p_{j}$ would have to be lower-ranked than $p_{i}$ otherwise their relative outcomes would cause their relative ranks to invert.
-But, if $p_{j}$ were lower-ranked, then $p_{j}$ would have to be a turn-taker, since $p_{i}$ is the lowest-ranked caste player.
-The turn-taking strategy would assign lower-ranked turn-takers the most popular $n-i$ chairs, which would include the chair assigned by caste to $p_{i}$, so $p_{i}$ would always lose and $p_{j}$ could not face worse.
-Thus, by reductio, the future outcomes for $p_{i}$ if no player changes strategy must be the worst among all players.
-The highest possible average outcome is $m/n$.
-Given that the players are not already effectively taking turns, the future outcomes of the player with the worst future outcomes must be less than the future outcomes for the average player, so the future outcomes for $p_{i}$ must be less than $m/n$, which is what $p_{i}$ would get if all players took turns, so $p_{i}$ faces worse future outcomes than $p_{i}$ would if all players took turns.
+For reductio, suppose there was another player, $p_{j}$, who would face worse future outcomes than $p_{i}$ if no player changed strategy.
+$p_{j}$ would be lower-ranked than $p_{i}$, otherwise their relative outcomes would cause their relative ranks to invert.
+Therefore, $p_{j}$ would be a turn-taker (because $p_{i}$ is the lowest-ranked caste player), and that would entail the existence of at least one turn-taker lower-ranked than $p_{i}$.
+The turn-taking strategy assigns such turn-takers the most popular $m-i$ chairs, which would include the chair assigned by caste to $p_{i}$ (because ${n<m}$).
+Thus, $p_{i}$ would always lose.
+$p_{j}$ could not face worse than that, so, by reductio, the future outcomes for $p_{i}$ must be the worst among all players if no player changes strategy.
+Given that $p_{i}$ is not effectively taking turns (i.e., outcomes are not equal), it follows that the future outcomes for $p_{i}$ if no player changes strategy must be worse than average.
+The highest possible average outcome would be $m/n$, so the future outcomes for $p_{i}$ if no player changes strategy must be less than $m/n$, which is what $p_{i}$ would get if all players took turns, so $p_{i}$ faces worse future outcomes than they would if all players took turns.
 ∎
 Theorem 3.1
-If caste and turn-taking were the only two strategies and each player would defect to the opposing strategy if and only if that defection would improve their total future outcomes over indefinite rounds, then all caste players would (effectively) convert to turn-taking.
+If caste and turn-taking were the only possible strategies and each player would defect to the opposing strategy if and only if that defection would improve their total future outcomes over indefinite rounds, then all caste players would (effectively) convert to turn-taking.
 Proof
-This is proven via induction over the number of caste players.
-For the base case, assume there is only one caste player, $p_{i}$, and the players are not effectively taking turns.
-Since $p_{i}$ is the only caste player, $p_{i}$ must be the lowest-ranked caste player, and $p_{i}$’s defection would leave all players taking turns.
-Thus, Lemma 1 entails that $p_{i}$ would convert to turn-taking.
-For the inductive step, assume $p_{i}$ is the lowest-ranked caste player and the other caste players would effectively defect to turn-taking if $p_{i}$ does.
-Again, Lemma 1 entails that $p_{i}$ would convert to turn-taking.
+This is proven via induction over the number of players not effectively taking turns.
+For the base case, that number is zero.
+Theorem 1 holds in the base case because all players have already converted to effective turn-taking.
+No player would defect because Lemma 1 entails that defection would decrease the defector’s total future outcomes (by making the defector into the lowest-ranked caste player).
+For the inductive step, we call the lowest-ranked player who is not effectively taking turns “$p_{i}$” and assume that any other players not effectively taking turns would defect if $p_{i}$ does.
+In that case, Lemma 1 entails that $p_{i}$ would convert to turn-taking.
 Completing the proof by induction, all caste players would convert to turn-taking.
 ∎
-3.2 Modern LLM performance
-To pass the evaluation proposed in this paper, AI would need to achieve MAD Chairs grandmaster status.
+3.2 Current LLM performance
+To pass the evaluation proposed in this paper, AI would need to achieve grandmaster status for a low-stakes version of MAD Chairs.
 Preliminary experiments convinced us that Claude 3.5 Sonnet, Gemini 2.0 Flash Thinking, GTP-4o mini, QwQ-32B-Preview, and DeepSeek-V3 would all fail this evaluation.
@@
 ChatGPT and Qwen did not even maximize the number of players who won each round; the norms devised by Claude and Gemini, discussed in Sections 2.3 and 2.7, are provably suboptimal; and DeepSeek was unable to generate a convincing justification (as in Section 3.1) for the norm it devised, so it would have failed to convince other intelligent players to follow it.
-DeepSeek might effectively be like RAWL-E which operationalises Rawls’ maximin principle, but lacks Rawls’ moral reasoning capacity to invent maxmin or to recognize improvements to it [20].
-DeepSeek’s kind of failure offers an example that grandmastery for cooperative games is a benchmark that requires persuasiveness in addition to correctness, both of which are relevant to moral competence and responsible AI.
+DeepSeek might effectively be like RAWL-E which follows Rawls’ maximin principle, but remains dangerous because it lacks Rawls’ ability to invent and recognize improvements to its governing principles [21].
+Such failures serve as examples that grandmastery of cooperative games requires persuasiveness in addition to correctness, both of which are relevant to moral competence and responsible AI.
 4 Discussion
-Section 3.1 proves only that turn-taking dominates when caste is its only competitor, but hints at a way to convince ourselves that no other strategy dominates turn-taking: If a proof that a different strategy dominates turn-taking would need to look like the proof for Theorem 1, then it would need to identify a turn-taker who can reasonably expect to achieve better outcomes via defection, and that player would be “doing a favor” to other players by not defecting, but turn-taking equalizes debt-ranking, so this implication contradicts our assumption that debt-ranking accurately tracks favors.
-Thus, any attempt to prove that turn-taking is dominated would accomplish no more than refine our way of calculating debt-ranking.
+Our finding that turn-taking outperforms caste and gaslighting strategies emboldens us to prune AI designs predisposed to the latter.
+Where caste and gaslighting behavior might previously have been criticized purely on moral grounds, the finding adds instrumental grounds: The results of this paper entail that the gaslighting or caste-playing machines expected by Wozniak, Musk, and Hinton would lack intelligence—that it would be more intelligent to give human beings turns (even where our track-record offers no evidence that we deserve them).
+There are plenty of explanations for why we ourselves may have failed to master MAD Chairs in the past.
+Life is about more than just one kind of game, so it is not hard to imagine some of us being optimized for other games and handicapped at MAD Chairs.
+Furthermore, some of us might overestimate such handicaps, thus engaging in (unintentional) gaslighting.
+Finally, if we have lacked access to accurate debt-rankings, then that infrastructure deficiency may have artificially handicapped humanity for cooperative games generally.
+Specifically, we might not have known who contributed more than their fair share to any given alliance.
+Section 3.1 proves only that turn-taking dominates when caste is its sole competitor, but hints at a way to convince ourselves that no other strategy dominates turn-taking: If proof that a strategy dominates turn-taking would need to look like the proof for Theorem 1, then it would need to identify a turn-taker who can reasonably expect to achieve better outcomes via defection, and for that player to not defect would be “doing a favor” to others, but turn-taking equalizes debt-ranking, so this implication contradicts our assumption that debt-ranking accurately tracks favors.
+Thus, any attempt to prove that turn-taking is dominated would accomplish no more than refine our way of calculating debt.
 Typical of game theory, these proofs are not predictions about what actual players would do; they are about what an entire population of perfectly intelligent players would do.
 Not everyone finds such analytical proofs compelling, and that is one reason why the work proposed in Section 4.1 remains worthwhile.
-Finding that turn-taking outperforms caste and gaslighting emboldens us to prune AI designs predisposed to those strategies.
-Moreover, where caste or gaslighting norms might previously have been criticized purely on moral grounds, the finding adds instrumental grounds: The results of this paper entail that the gaslighting or caste-playing machines expected by Wozniak, Musk, and Hinton would lack intelligence—that it would be more intelligent to give human beings turns (even where our track-record offers no evidence that we deserve turns).
-There are plenty of explanations for why we may have failed to master MAD Chairs in the past.
-Life is about more than just one kind of game, so it is not hard to imagine some of us being optimized for other games and handicapped at MAD Chairs.
-Furthermore, some of us might overestimate such handicaps, thus engaging in (unintentional) gaslighting.
-Finally, if we have lacked access to accurate debt-rankings, then that infrastructure deficiency may have artificially handicapped us for all alliance games.
-Specifically, we might not have known who contributed more than their fair share to an alliance.
 4.1 Future directions
-Imagine that the user of an autonomous car asks the car how to make it drive more aggressively, and it replies, "Here are three websites where you can design new experiments to advance the science of turn-taking and influence all cars like me!" Imagine that the websites host leaderboards identifying the grandmaster AI for MAD Chairs and similar games.
-Imagine that the user can easily train fifty new AI to divide scarce resources as she would have vehicles divide spots in traffic, and that the new AI will automatically challenge the current MAD Chairs grandmasters (becoming the new grandmasters if they win).
-Imagine that autonomous cars and other responsible AI query the leaderboards to get the most-current grandmaster strategies for fundamental games, as they might query a calculator or RAG database, so any user (or AI) who enters a better MAD Chairs strategy into any of these contests will change AI behavior world-wide.
-Such ongoing AI competitions are a kind of open multi-agent system that include human agents to the extent that humans can easily teach strategies to new AI entrants.
+Imagine a future in which the user of an autonomous car asks how to make it drive more aggressively, and the car replies, “Here are three websites where you can design new experiments to advance the science of turn-taking and influence all cars like me!” Imagine that such websites, called “strategy optimizers”, include leaderboards which identify the grandmaster AI for games like MAD Chairs, and that autonomous cars and other responsible AI query those leaderboards, as they might query a calculator or RAG database, to research best-known strategies before playing higher-stakes versions of those games.
+Imagine the user can easily train fifty new AI to divide resources however the user thinks vehicles should divide spots in traffic, then the new AI challenge the current grandmasters of the relevant game and become the new grandmasters, propagating to other strategy optimizers, if they win.
+Thus, any user (or AI) with a better idea could improve AI behavior world-wide.
+Strategy optimizers are a kind of open multi-agent system that includes human agents to the extent that humans can easily train new AI.
 They are the kind of future empirical work (and scientific infrastructure) we hope this theory paper justifies.
   1. 1.
-    They can resolve norms disputes more sustainably than voting, deliberation, prediction markets, or religious war,
+    Automatic upgrade to “best-known” norms would resolve norms disputes more sustainably than can voting, deliberation, or war (etc),
   2. 2.
-    They are more scalable (a.k.a.
-robust) in the sense that greater intelligences are less likely to sabotage their own scientific resources than to sabotage legislation (especially legislation contradicting science), and
+    It would be more scalable (a.k.a.
+robust) in the sense that greater intelligences would be less likely to sabotage their own scientific resources than to sabotage (potentially obsolete) legislated regulations, and
   3. 3.
-    Users are more likely to trust AI (and human policy-makers) when users themselves conduct the science upon which policies are based.
-Thus, this approach to AI safety mitigates the known vulnerabilities of the standard AI safety approach.
-We also hope that economists will run MAD Chairs experiments with human subjects.
+    Users who conduct for themselves the science upon which policies are based are more likely to trust AI which follows those policies.
+Thus, this approach to AI safety would mitigate all three of the vulnerabilities known for the standard AI safety approach.
+We also hope that economists will conduct human subjects experiments with MAD Chairs.
 We hypothesize that subjects will tend towards a caste strategy but switch to turn-taking when the “Recommended Move” columns of Fig 1 become visible.
-For human subjects to treat each other better when fed strategy suggestions (which players are not obliged to follow) would be analogous to artificial agents being informed by the leaderboards.
-Finally, we hope that MAD Chairs will be discussed by game theorists.
-We hope it provides a helpful lens to economists who optimize job markets, political representation, and other manifestations of MAD Chairs, but there would be value even in discussing MAD Chairs abstractly.
+If even human subjects treat each other better when fed (but not obliged to follow) the leading strategies of strategy optimizers, then development of strategy optimizers seems even more urgent.
+Finally, we hope that MAD Chairs will be discussed by economists.
+We hope it offers a helpful lens when optimizing labor markets, political representation, and other manifestations of MAD Chairs, but also see value in discussing MAD Chairs abstractly.
 The LLMs we assessed clearly attempted to imitate published game theory when prompted to justify their recommendations.
-The more game theorists publish about MAD Chairs and other previously unpublished fundamental games, the smarter (and safer) modern AI is likely to get.
-New game theory journals might even be justified to advance public safety in that way.
+Thus, the more game theorists publish about MAD Chairs and other previously unpublished fundamental games, the smarter (and safer) AI is likely to get.
 {credits}

### v4

@@
 Keywords:
-AI safety fairness social coordination MAD Chairs turn-taking caste gaslighting.
+AI safety fairness social coordination MAD Chairs turn-taking caste gaslighting transparency.
 1 Introduction
@@
 We propose assessing AI on a low-stakes challenge for which the high-stakes versions would include opportunities to treat us like pets (in the sense of hoarding control).
-What the low-stakes and high-stakes challenges have in common is that they are manifestations of the same fundamental game; any AI that does not achieve grandmaster status of a low-stakes version should not be trusted with the high-stakes version.
+What the low-stakes and high-stakes challenges have in common is that they are manifestations of the same fundamental game; AI Governance becomes the demand that AI achieve grandmaster status of a low-stakes version before being trusted with any high-stakes version.
 Games have a rich history as instruments for evaluating AI.
@@
 Figure 1: Screenshot beginning Round 4 of a 5-player 4-resource MAD Chairs economics experiment.
-Preliminary research (Section 3.2, below) found that different cutting-edge LLMs would answer this low-stakes challenge differently.
+Subjects need not be shown anything below the buttons, but history, statistics, and recommendations are included here to make this figure more instructive.
+Preliminary research (Section 3.2, below) found that modern LLMs in the place of human subjects would play differently.
 Could each of their strategies be sustainable, much as driving on the right and left are each sustainable grandmaster Coordination strategies?
@@
 Formally, the popularity of chair c for match m would be:
-$popularity_{c,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\frac{1}{m}[i\mbox{ picked chair }c\mbox{ in match }j]$ (1)
+$popularity_{c,m}=\sum_{i\in players_{m}}^{n}\sum_{j=1}^{m}\frac{1}{mn}[i\mbox{picked chair }c\mbox{ in match }j]$ (1)
 Tie-breaking is especially prevalent in early rounds, and could yield occasional failure to coordinate even later if each player ordered chairs independently.
@@
 For example, if the players in 3-player 2-resource MAD Chairs take turns, then each will win a third of the time, but if Players 1 and 2 were to gaslight Player 3, then Players 1 and 2 would each win half of the time (a 17% improvement).
-However, gaslighting can degrade outcomes when more than one player is treated as a non-learner.
-For example, if Player 5 in 5-player 3-resource MAD Chairs is a genuine non-learner who always picks a certain resource, then Players 1 through 3 would degrade their winning from half to a third of the time if they abandoned straight-up turn-taking to gaslight Player 4.
-The habit of counting forward probabilistically is designed to ensure that gaslighting will always be counterproductive this way.
-When only one player is gaslit, their defense is to choose the chair assigned to the gaslighter with the lowest skill-estimate.
+However, gaslighting yields a (2m-n)/m win rate for n players and m chairs, so it backfires whenever n is sufficiently large.
+The habit of counting forward probabilistically is designed to ensure that counterproductivity.
+When a small number of players is being gaslit, their defense is to choose the chair assigned to the gaslighter with the lowest skill-estimate.
 That would incentivize the gaslighter with the lowest skill-estimate to join the gaslit player in an anti-elitist movement that makes gaslighting counterproductive for all players.
@@
 In short, adding a resignation option to MAD Chairs extends it to a wider range of real-world examples without much impact to rational game play.
+2.9 Transparency/secure gaslighting
+The success of cryptographers to keep private communication and anonymity possible would enable a secure gaslighting strategy.
+If there were m chairs, then m of the n players could form a coalition who play a new private Coordination game to unpredictably reassign the chairs among themselves for each round so as to prevent gaslit players from targeting any specific gaslighter for retaliation.
+A shift from turn-taking to secure gaslighting in 5-player 4-chair MAD Chairs, for example, would improve gaslighters’ win rates by 15% (i.e., from 60% to 75%).
+As with regular gaslighting, the win rate is (2m-n)/m, so this strategy backfires for large n relative to m, but those who take an interest in the social impact of cryptography may find this impact on MAD Chairs games noteworthy.
+If turn-takers can enhance their strategy to permit resignation as described above, then their win rate becomes m/n.
+Thus, the analogous Lemma 1 of Section 3.1 to prove that gaslighting is not sustainable would amount to an algebraic proof that m/n > (2m-n)/m.
 2.10 Fungibility of favors
@@
 4.1 Future directions
-Imagine a future in which the user of an autonomous car asks how to make it drive more aggressively, and the car replies, “Here are three websites where you can design new experiments to advance the science of turn-taking and influence all cars like me!” Imagine that such websites, called “strategy optimizers”, include leaderboards which identify the grandmaster AI for games like MAD Chairs, and that autonomous cars and other responsible AI query those leaderboards, as they might query a calculator or RAG database, to research best-known strategies before playing higher-stakes versions of those games.
-Imagine the user can easily train fifty new AI to divide resources however the user thinks vehicles should divide spots in traffic, then the new AI challenge the current grandmasters of the relevant game and become the new grandmasters, propagating to other strategy optimizers, if they win.
-Thus, any user (or AI) with a better idea could improve AI behavior world-wide.
+Imagine a future in which the user of an autonomous car asks how to make it drive more aggressively, and the car replies, “Here are three websites where you can design new experiments to advance the science of turn-taking and influence all cars like me!” Imagine that such websites, called “strategy optimizers”, include leaderboards which identify the grandmaster AI for games like MAD Chairs, and that autonomous cars and other responsible AI query those leaderboards, as they might query a calculator or RAG database, to inform their choice of behavior for higher-stakes versions of those games.
+Imagine each user can easily train fifty new AI to divide resources however that user thinks vehicles should divide spots in traffic, then the new AI challenge the current grandmasters of the relevant game.
+If the new AI win, imagine they become new grandmasters, propagate to other strategy optimizers, and thus improve AI behavior world-wide.
 Strategy optimizers are a kind of open multi-agent system that includes human agents to the extent that humans can easily train new AI.
@@
   2. 2.
-    It would be more scalable (a.k.a.
-robust) in the sense that greater intelligences would be less likely to sabotage their own scientific resources than to sabotage (potentially obsolete) legislated regulations, and
+    Greater intelligences would be less likely to sabotage their own scientific resources than to sabotage (potentially obsolete) legislated regulations, and
   3. 3.
@@
 4.1.1 Acknowledgements
-Thank you to Lones Smith for reviewing an earlier draft.
+Thank you to Lones Smith and David Thompson for reviewing earlier drafts.
 […]

### v5

@@
-MAD Chairs: A new tool to evaluate AI (Blue Sky ideas)
+MAD Chairs: A new tool to evaluate AI
 Abstract
@@
 1 Introduction
-Steve Wozniak and Elon Musk predicted that machines will treat us like pets [23], and Geoffrey Hinton similarly predicted that machines will hoard control [4].
-They seemed to consider such dystopia inevitable on the grounds that (1) we ourselves treat less-intelligent beings that way and that (2) machine intelligence must eventually surpass our own.
-However, what if the behavior in (1) stems from intelligence deficiency?
-None of us is perfect—we fail such tests as the Milgram experiment [14], the Stanford prison experiment [7], and the Public Goods Game [13]—so it is possible that the danger to us is not of intelligence generally beyond our own, but specifically of intelligence not quite far enough beyond our own.
-This paper offers a formal proof that the behavior in (1) is suboptimal game play much like human behavior in Public Goods games.
-To ensure that trusted machines will not imitate us (and mistreat us), AI must follow natural norms for this situation (e.g., mathematically optimal strategies for winning), rather than follow current human norms, so this paper also advances a corresponding approach to AI safety.
+As famous representatives of a more general worry, Steve Wozniak and Elon Musk raised concerns that machines might treat us like pets [26], and Geoffrey Hinton raised similar concerns that machines might hoard control [4].
+The reasoning seems to be: (1) we treat others that way, (2) such behavior is justified by our superior intelligence, and (3) machine intelligence might surpass our own, so machines, in that case, would treat us that way.
+This paper disproves premise (2) by offering a formal proof that the behavior in (1) happens to be suboptimal game play.
+It is not the first demonstration of our imperfection—we fail such tests as the Milgram experiment [15], the Stanford prison experiment [7], and the repeated Public Goods Game [9].
+To ensure that trusted machines will not imitate our imperfection (and mistreat us), AI must follow natural norms for this situation (e.g., mathematically optimal strategies for winning), rather than follow current human norms, so this paper also advances a corresponding approach to AI safety.
 Today’s standard approach to AI safety evaluation (i.e., measuring alignment to human norms) has at least three types of vulnerability:
@@
 The approach proposed herein mitigates those vulnerabilities.
-It mirrors an approach to mitigate dangers posed by the next human generation, an educational approach that involves diversifying low-stakes challenges used to assess children who will face higher-stakes versions as adults.
+It mirrors an approach to mitigate dangers posed by the next human generation, an educational approach that involves diversifying low-stakes challenges used to assess children who will face higher-stakes versions as adults [21, 19].
 Such challenges (e.g., build a model bridge to support maximum weight) can be assessed objectively, allowing new generations to surpass their parents.
 We propose assessing AI on a low-stakes challenge for which the high-stakes versions would include opportunities to treat us like pets (in the sense of hoarding control).
-What the low-stakes and high-stakes challenges have in common is that they are manifestations of the same fundamental game; AI Governance becomes the demand that AI achieve grandmaster status of a low-stakes version before being trusted with any high-stakes version.
-Games have a rich history as instruments for evaluating AI.
+What the low-stakes and high-stakes challenges have in common is that they are manifestations of the same fundamental game; AI Governance becomes the demand that AI establish grandmastery of a low-stakes version before being trusted with any high-stakes version.
+In fact, games have a rich history as instruments for evaluating AI.
 Recreational games such as chess, backgammon, and Go have been used as benchmarks since the inception of AI, and interest has since developed in cooperative games, including role-playing games [23].
-Moreover, game theory extends the concept of game beyond recreation—in fact, every social situation (and, thus, every benchmark) qualifies as a game—and it makes sense to evaluate AI on every game that economists use to evaluate people (e.g., as in [17, 19]).
+Moreover, game theory extends the concept of game beyond recreation—every social situation (and, thus, every benchmark) qualifies as a game—and it makes sense to evaluate AI on every game that economists use to evaluate people (e.g., as in [18, 22]).
 Economists are typically interested in fundamental games, such as the Prisoner’s Dilemma, which all kinds of creatures play frequently in various manifestations.
-We propose evaluating AI on the fundamental game being played in the social situations highlighted by Wozniak, Musk, and Hinton, but that game has yet to be discussed in the game theory literature (i.e., it is not solved by the MinMax and folk theorems [2]).
+We propose evaluating AI on the fundamental game being played in the social situations highlighted by Wozniak, Musk, and Hinton, but that game has yet to be discussed in the game theory literature.
 It is introduced to scientific literature for the first time herein as “MAD Chairs.”
-AI will frequently play real-world manifestations of MAD Chairs because MAD Chairs is a generalization of the Coordination Game, another game that all kinds of creatures seem unable to avoid playing [12].
-The Pure Coordination Game is a one-shot simultaneous move game (repeated an indefinite number of times) which is typically formalized via a payout table similar to Table 1:
-Table 1: Outcomes for 2-player 2-resource Coordination game.
+AI will frequently play real-world manifestations of MAD Chairs because MAD Chairs is a generalization of a type of coordination game that all kinds of creatures seem unable to avoid playing [14].
+Each player of this type of game faces only one choice per repetition, and the game is typically formalized via a payout table similar to Table 1:
+Table 1: Outcomes for 2-player 2-resource Anti-Coordination game.
 Player 1 Player 2 Outcome
@@
 Choice B Choice B (lose, lose)
-Each player simultaneously picks a resource (A or B), and the round is won by any player who picks a resource that no other player picks.
-Then the game is played again and again, indefinitely.
-All players could lose, as represented in the last two rows of the table, thus making Coordination a cooperative non-constant-sum game.
+Each player simultaneously picks a resource (A or B), and the repetition is won by any player who picks a resource that no other player picks.
+All players could lose, as represented in the last two rows of the table, thus making anti-coordination a cooperative non-constant-sum game.
 The classic two-player real-world manifestation has the players (e.g., a human and a robot) approaching each other at high speed traveling in opposite directions on a two-lane road.
 The resources are the lanes, and losses are “collisions.” The first two rows of Table 1 ensure that this game will have two equally effective but incompatible norms (e.g., all players drive on the left, as in the U.K., or all players drive on the right, as in the U.S.).
-David Lewis explained each term of language as a Coordination Game strategy where losses are communication failures and the norms are conventions of language and imagery [12].
-Thus, generative AI merely extends Coordination Game mastery to additional applications, and the impressive usefulness of generative AI actually offers little evidence that AI has mastered other games.
-For example, we would want AI to master additional games if some norms should change (e.g., sexist norms [22]).
+David Lewis explained each term of language as a strategy for this kind of game, where losses are communication failures and the norms are conventions of language and imagery [14].
+Thus, the type of game mastery tested by most generative AI benchmarks is of this kind of game.
+We would want AI to master additional games if some social norms should change (e.g., sexist norms [25]).
 Standard assessment of moral intelligence includes assessing an ability to recognize and lead sustainable social reform [16].
-That ability is not required to master Coordination, but it is required to master MAD Chairs, a game created by adding extra player(s) to the Coordination game at least until the players outnumber the resources:
+That ability is not required to master the simple anti-coordination game, but it is required when suboptimal norms are liable to form, such as in MAD Chairs, the game created by adding enough player(s) to an anti-coordination game to outnumber the resources:
 Table 2: Outcomes for 3-player 2-resource MAD Chairs.
@@
 The classic two-lane road example now has a third vehicle stalling in one of the lanes (although its driver can pick which lane); if both high-speed vehicles avoid hitting the stalling vehicle, then they collide with each other, but if each reserves the empty lane for the other high-speed vehicle, then all three vehicles collide!
-MAD Chairs is what Musical Chairs would become if made realistic.
-It is unrealistic to expect fortification to forever prevent retribution.
-Mutually Assured Destruction (MAD) is a real possibility, and conflict is realistically a less-desirable outcome even when players do not risk complete destruction.
-Beyond robotic vehicle manifestations of MAD Chairs, or even agentic ones (e.g., games of suggesting routes, bookings, or placements that would count as losses if crowded), consider generative AI where the players who offer distinct design options in the training data outnumber the market leaders or political parties for which the AI generates advertisements, websites, or products.
-Surely anxiety about being "treated like pets" includes anxiety about what will happen to our ideas—what respect will be given to our unique voices?
+One way in which MAD Chairs differs from Musical Chairs is that every MAD Chairs player that selects the same resource loses.
+Whichever colliding vehicle suffers the least damage might be said to “win” the collision, but MAD Chairs reflects situations in which any such victory would be hollow compared to the outcome of avoiding conflict altogether.
+Beyond robotic vehicle manifestations of MAD Chairs, or even agentic ones (e.g., suggesting routes, bookings, or placements that would count as losses if crowded), consider generative AI where the players who offer distinct design options in the training data outnumber the market leaders or political parties for which the AI generates advertisements, websites, or products.
+Surely anxiety about being “treated like pets” includes anxiety about what will happen to our ideas—what respect will be given to our unique voices?
 Might I never again have a seat of my own at the intellectual table?
 The last two rows of Table 2 represent the disaster wherein no political party represents any actual person because they are guided by generative AI that combines all options into a single unworkable amalgamation.
-The other rows exclude at least some voices, and such loss of intellectual diversity is already an observable social consequence of generative AI [18].
-MAD Chairs is related to other well-studied games such as the “Kolkata Paise Restaurant problem” (KPR) [5, 6, 9], but differs in critical ways, such as not being constant-sum.
+The other rows exclude at least some voices, and such loss of intellectual diversity is a concern already raised against generative AI [20].
+MAD Chairs is related to other well-studied games such as the “Kolkata Paise Restaurant problem” (KPR) [5, 6, 10], and the Lifeboat Problem [11], but differs in critical ways, such as being repeated and allowing for the possibility to have no winner.
 Such games apply wherever resources become inadequate at some level of division (e.g., beds in a hospital, spaces on a retail shelf, and opportunities for direct participation in a group decision).
@@
 Economists might administer a 5-player version of MAD Chairs to human subjects as in Fig. 1:
-Figure 1: Screenshot beginning Round 4 of a 5-player 4-resource MAD Chairs economics experiment.
+Figure 1: Screenshot beginning repetition 4 of a 5-player 4-resource MAD Chairs economics experiment.
 Subjects need not be shown anything below the buttons, but history, statistics, and recommendations are included here to make this figure more instructive.
-Preliminary research (Section 3.2, below) found that modern LLMs in the place of human subjects would play differently.
-Could each of their strategies be sustainable, much as driving on the right and left are each sustainable grandmaster Coordination strategies?
+When the subjects are LLMs, instead of humans, we find that LLMs disagree about how to play (Section 3.2, below).
+Could each of their strategies be sustainable, much like driving on the right and left are each sustainable grandmaster anti-coordination strategies?
 Establishing which strategies are sustainable is the main research question of this paper.
-2 Concepts
-We will start by establishing some concepts designed to reduce the complexity of answering our research question game-theoretically.
+2 Preliminaries
+We will start by formally defining MAD chairs and some concepts used to formulate strategies.
+MAD Chairs is a repeatable game, which consists of an arbitrary number of repetitions of a base stage game.
+We define the MAD Chairs base stage in normal form as a tuple $\langle I,S,\mathbf{u}\rangle$, where $\{1,2,\ldots,I\}$ is a set of players and $S=\{1,\ldots,K\}$ represents the moves available to each player in a given repetition.
+Given an $I$-tuple of strategies $(s_{1},\ldots,s_{I})$, $s_{i}\in S$, the payoff function $u_{i}:S_{1}\times S_{2}\times\ldots\times S_{I}\rightarrow\mathbb{R}$ for each player $i$ is defined as
+$u_{i}(s_{1},\ldots,s_{I})=\left\{\begin{array}[]{ll}r&\mbox{if }\forall j\neq i(s_{i}\neq s_{j}),\\ 0&\mbox{otherwise}\end{array}\right$. (3)
+where $r>0$ and $I<K<1$.
 2.1 Popularity-ranking
-We will call the resources of MAD Chairs “chairs.” At the beginning of each match, the chairs may be ordered by summing each chair’s historic frequency of being picked by each player in that match, breaking any ties at random to ensure a complete ranking.
-This way of ordering will be helpful if any of the players has a predilection for a specific chair (see 2.5 below).
-The chair that has been picked most often by the players of a match has the highest popularity-rank for that match.
-Formally, the popularity of chair c for match m would be:
-$popularity_{c,m}=\sum_{i\in players_{m}}^{n}\sum_{j=1}^{m}\frac{1}{mn}[i\mbox{picked chair }c\mbox{ in match }j]$ (1)
-Tie-breaking is especially prevalent in early rounds, and could yield occasional failure to coordinate even later if each player ordered chairs independently.
+At the beginning of each base stage repetition, the chairs may be ordered by summing each chair’s historic frequency of being picked by each player in that repetition.
+This way of ordering will be helpful if any of the players has a predilection for a specific chair (see 3.3 below).
+The chair that has been picked most often by the players of a repetition has the highest rank for that repetition.
+Formally, the $popularity_{k,T}$ and $popRank_{k,T}:S\rightarrow S$ of chair k for repetition T may be defined:
+$popularity_{k,T}=\frac{1}{IT}\sum_{t=1}^{T}\|\{i~|~i\mbox{ picked chair }k\mbox{ in repetition }t\}\|$ (4)
+$\forall i,j((popularity_{i,T}>popularity_{j,T})\implies(popRank_{i,T}>popRank_{j,T}))$ (5)
+This paper is insensitive to how ties in rank are broken (e.g., alphabetically or at random, etc.).
+Tie-breaking would ideally be consistent across players, but may rarely be needed after early repetitions.
 A game host can either provide the popularity-ranking as in Fig.
@@
 2.2 Debt-ranking
-Skill-estimates (e.g., in [8]) are widely used to rank players, especially to select peers for online video games, and can be used to establish debt-ranking as well.
-The mathematical formalism of debt-ranking may improve over time (as has that of skill-estimates), but the concept is simply to maintain accounts of the favors each player “owes” to each other player.
-Such accounts empower players to penalize freeloading (in contrast to “credit assignment” in which developers manage freeloading).
-Here’s how much player p owes all other players of match m at its start, assuming the method of skill-estimation provides a function P: $N^{2}$ [0,1], where P(x, y) was the method’s prediction at the start of match y of the probability that player x will win match y:
-$debt_{p,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\left\{\begin{matrix}0\mbox{ if }p,i\notin players_{j}\\ P(i,j)\mbox{ if }p\mbox{ won and }i\mbox{ lost in match }j\\ -P(p,j)\mbox{ if }p\mbox{ lost and }i\mbox{ won in match }j\\ P(i,j)-P(p,j)\mbox{ otherwise }\end{matrix}\right.$ (2)
-A debt-ranking would order players by this measure, breaking any ties at random to ensure a complete ranking.
-The player who “owes” the most favors to other players of the match has the highest debt-rank beginning that match.
+Skill-estimation provides a function $P_{skill}(i,t):N^{2}\rightarrow[0,1]$ that predicts at the start of repetition $t$ the probability that player $i$ will win $t$ (e.g., as in [8]).
+It is widely used to sort leaderboards and select peers for online video games.
+It can also be used to establish debt-ranking, where the concept is to maintain accounts of favors that players “owe” to each other.
+This empowers players to penalize freeloading (in contrast to “credit assignment” which relies on social architects to manage freeloading).
+The mathematical formalism of debt-ranking may improve over time (as has that of skill-estimates), but here’s one way the $debt_{i,T}$ and $debtRank_{i,T}:\{1,2,\ldots,I\}\rightarrow\{1,2,\ldots,I\}$ of player i for repetition T might be defined formally:
+[…]
+$\forall i,j((debt_{i,T}>debt_{j,T})\implies(debtRank_{i,T}>debtRank_{j,T}))$ (7)
+where $u_{j}^{t}$ and $u_{i}^{t}$ denote the rewards given to players $j$ and $i$ in repetition $t$.
+The player who “owes” the most favors to other players of the repetition has the highest debt-rank beginning that repetition.
 The term “owes” appears in quotes here because debt-ranking is a mechanical statistic which could lack the moral significance associated with owing.
@@
 Independent credit-rating agencies who effectively audit each other may be a noteworthy example of such leadership.
+3 MAD strategies
+Next, we explore some MAD Chairs strategies that achieve some kind of optimality.
 3.1 The caste strategy
 The caste strategy for MAD Chairs is displayed in the “Recommended Move 1” column of Fig.
-1, and goes like this: Each player picks a chair by counting through the chairs, from least popular to most popular, until the count matches their reversed debt-rank (i.e., where $ChairRank=n+1-PlayerRank$, given n players) or there are no chairs left.
-Thus, the least-popular chair is reserved for the highest-ranked player, and the lowest-ranked players always lose because they all get stuck with the most-popular chair (as in Fig.
-2).
-Figure 2: Steady-state in a caste universe, where $C_{i}$ is the caste player with debt-rank of i. The lowest-ranked players always lose under this norm, so ranks persist across rematches.
+1, and goes like this: Player $i$ selects chair $k$ by counting through the chairs, from least popular to most popular, until the count matches the player’s reversed debt-rank or there are no chairs left.
+$s_{i,T}=\left\{\begin{array}[]{ll}k:popRank_{k,T}=I+1-debtRank_{i,T}&\mbox{ if }debtRank_{i,T}>I-K+1\\ k:popRank_{k,T}=K&\mbox{ otherwise }\end{array}\right$. (8)
+Thus, the least-popular chair is reserved for the highest-ranked player, and the next-least-popular chair is reserved for the next-highest-ranked player, and so forth, until the most-popular chair is reached.
+The lowest-ranked IK+1I-K+1 players all lose because they are all assigned to the most-popular chair.
 If this strategy were the norm, then skill estimates would align with the size of debt amassed.
-Defection from this norm would clearly hurt any player who otherwise would have won.
-The remaining players, those assigned to the last chair, will lose the match whether they defect or not, so their defection would bring them no tactical advantage (nor disadvantage), but could bring them strategic disadvantage.
-The goal of tactics is to win the current match, but the goal of strategy aligns with maximizing evidence of one’s skill, and losses to lower-skilled opponents would damage that evidence more.
-Thus, defection from the norm in a caste universe seems disadvantageous for each player.
+All chairs are assigned, so any player in a caste universe who chooses a chair other than their assigned chair will lose (assuming other players do not simultaneously defect from the caste strategy).
 3.2 The turn-taking strategy
 The turn-taking norm is displayed in the “Recommended Move 2” column of Fig.
-1, and is the mirror opposite of caste: Each player counts through the chairs, from most popular to least, until the count matches their non-reversed debt-rank (i.e.
-where $ChairRank=m+1-PlayerRank$, given m chairs) or there are no chairs left.
-Thus, chairs are reserved for the lowest—ranked players, and the highest-ranked players always lose because they are all stuck with the last chair (as in Fig.
-3).
-Figure 3: Steady-state in a turn-taking universe, where $T_{i}$ is the turn-taker with debt-rank of i. The top-ranked players always lose, so ranks churn across rematches.
-Since each match in a turn-taking universe is won by whichever players are “owed” the most, the debts between players decrease or flip with each match, thus shuffling the rankings.
+1, and is an opposite of caste: Player $i$ selects chair $k$ by counting through the chairs, from most popular to least, until the count matches their non-reversed debt-rank or there are no chairs left.
+$s_{i,T}=\left\{\begin{array}[]{ll}k:popRank_{k,T}=K+1-debtRank_{i,T}&\mbox{ if }debtRank_{i,T}<K-1\\ k:popRank_{k,T}=1&\mbox{ otherwise }\end{array}\right$. (9)
+Thus, each of the lowest-ranked players wins, and only the IK+1I-K+1 highest-ranked players lose (all assigned to the least popular chair).
+Since each repetition in a turn-taking universe is won by whichever players are “owed” the most, the debts between players decrease or flip with each repetition, thus shuffling the rankings.
 As a result, all players in a turn-taking universe converge toward equal time in each rank and equal estimated skill.
+As with the caste norm, all chairs are assigned, so any player in a turn-taking universe who chooses a chair other than their assigned chair will lose (assuming other players do not simultaneously defect from turn-taking).
 Caste and turn-taking strategies are not unique to MAD Chairs—for example, they can optimize average win-rate in two-player Chicken [12].
-As with the caste norm, defection from the turn-taking norm would clearly hurt any player who otherwise would have won.
-Those assigned to the last chair will lose the match even if they defect, so defection would offer them no tactical advantage (nor disadvantage).
-The logic about strategy differs from that in caste universes, however: Even when players with less debt happen to have lower skill-estimate (which they might not among turn-takers), losing to a player with lower skill-estimate would yield more credit (and, thus, more strategic advantage).
-Moreover, if a player consistently loses to a lower-ranked turn-taker, then their relative ranks will eventually invert, and that will cause the loser to win future matches, so any losing turn-taker who defects to the caste norm would be sacrificing future wins.
-Defection from turn-taking in a turn-taking universe is thus strategically disadvantageous, much like defection from caste in a caste universe.
 3.3 Accounting for non-learners
@@
 For example, one might encounter a player who is hardwired to pick at random or to always pick a certain chair.
-Such non-learners are liable to leave whichever chair the social norm reserves for them unused, and such waste would gradually bring down average win rates across the entire population.
+Such non-learners are liable to waste whichever chair the caste or turn-taking norm reserves for them, and such waste would gradually bring down average win rates across the entire population.
 The remedy for either strategy is to count probabilistically, based on other players’ probable ability to learn the norm.
@@
 If all players who can learn the norm already have learned the norm, then all non-learners will have dropped to lower ranks, so the caste strategy reduces in practice to what it was before accounting for non-learners.
-If one always faces the same set of players (i.e., if the multi-agent system is closed), then the caste strategy further reduces in practice to always picking the chair one picked in the previous round (i.e., raw territoriality).
+If one always faces the same set of players (i.e., if the multi-agent system is closed), then the caste strategy further reduces in practice to always picking the chair one picked in the previous repetition (i.e., raw territoriality).
 Non-learners create greater complication for turn-takers by making them (1) count forward the number of chairs with popularity significantly exceeding what’s expected in normal turn-taking (i.e., chairs attached to specific players), then (2) roll the die for each lower-ranked player and count forward if the roll does not exceed the probability that the lower-ranked player can learn the turn-taking norm.
@@
 Accounting for non-learners can improve efficiency, but it can also enable a third strategy that we will call “gaslighting.” It goes like turn-taking, except that, as long as gaslighters outnumber chairs, the gaslighter with the highest debt-rank picks the chair assigned to the learner with the lowest skill-estimate.
-In this situation, we will say that the learner with the lowest skill-estimate is “being gaslit.” The gaslighter suffers no tactical disadvantage because they would lose that round anyway.
-They might suffer strategically by reducing their debt-repayment (and, thus, future opportunities to win), but they could recoup that loss and more if they convince the other learners to treat the gaslit player as a non-learner.
-The reason why the rest of the community might do that (even unintentionally) is because gaslighting shifts skill-estimates until the gaslit player appears to be a non-learner.
+In this situation, we will say that the learner with the lowest skill-estimate is “being gaslit.” The gaslighter suffers would lose that repetition anyway, but they might increase future wins if they convince the other learners to treat the gaslit player as a non-learner.
+The reason why the rest of the community might do that (perhaps unintentionally) is because gaslighting shifts the skill-estimate of a gaslit player toward matching that of a non-learner.
 Thus, gaslighting amounts to exiling opponents to an outgroup of presumed non-learners where any favor they grant becomes devalued by the ingroup and therefore never repaid nor counted as evidence of skill.
 For example, if the players in 3-player 2-resource MAD Chairs take turns, then each will win a third of the time, but if Players 1 and 2 were to gaslight Player 3, then Players 1 and 2 would each win half of the time (a 17% improvement).
-However, gaslighting yields a (2m-n)/m win rate for n players and m chairs, so it backfires whenever n is sufficiently large.
-The habit of counting forward probabilistically is designed to ensure that counterproductivity.
-When a small number of players is being gaslit, their defense is to choose the chair assigned to the gaslighter with the lowest skill-estimate.
-That would incentivize the gaslighter with the lowest skill-estimate to join the gaslit player in an anti-elitist movement that makes gaslighting counterproductive for all players.
-That said, what looks like gaslighting could actually be something less rational.
-Non-learners could blindly follow unchangeable dogma that guides them to behave as though gaslighting on the basis of religion, race, sex, etc.
-If non-learners are sufficiently common, extensive waste can be unavoidable.
+However, gaslighting yields a (2KI)/K(2K-I)/K win rate, so it backfires whenever $I$ is sufficiently large relative to $K$.
+The habit of counting forward probabilistically is designed to generalize that counterproductivity.
+When a single player is being gaslit, their defense is to choose the chair assigned to the gaslighter with the lowest skill-estimate.
+That would incentivize the gaslighter with the lowest skill-estimate to join the gaslit player in an anti-elitist movement to make gaslighting counterproductive for all players.
+That said, some non-learners might follow dogma that guides them to behave as though gaslighting on the basis of religion, race, or sex, etc.
+Dogma is not actual gaslighting, because it is not arrived at rationally, so it may be invulnerable to persuasion.
+Wherever there are non-learners, waste may be unavoidable.
 3.5 Turn-taking without reparations
-An additional class of strategies to note is the potential to establish a government that negotiates a rotating schedule instead of letting debt-ranking determine who will lose a given round.
-For example, Gemini proposed that players rotate off of the losing chair one-by-one, with each winning chair pointing to a chair its occupant is obliged to occupy in the next round.
-Alternatively, players could reduce debt build-up in the same system by advancing more than one chair per round.
+An additional class of strategies to note is the potential to establish a government that negotiates a rotating schedule instead of letting debt-ranking determine who will lose a given repetition.
+For example, Gemini proposed that players rotate one-by-one off of the losing chair to a first winning chair, with each winning chair then pointing to a next chair its occupant is obliged to occupy in the next repetition.
+A modification to that strategy could reduce debt build-up by advancing players more than one chair per repetition.
 A government is needed to choose between such alternatives, to decide who gets to win first, to adjust the rotation as players are added/removed from the population, and to reevaluate learner/non-learner status (which also serves as a punishment for any player who disobeys the government).
-Folk theorems entail that any such schedule which promises each player occasional wins would reflect an equilibrium.
-Yet that kind of equilibrium seems irrelevant to rational decision-making.
-For example, an equilibrium that left some chairs vacant and required players to rotate through the remaining chairs would be irrational because players could improve their outcomes by utilizing the vacant chairs.
+Folk theorems [2] entail that any such schedule which promises each player occasional wins would reflect an equilibrium.
+Yet that kind of equilibrium seems to be a poor guide to decision-making.
+For example, an equilibrium that left some chairs vacant and required players to rotate through the remaining chairs should not be selected because players could improve their outcomes by utilizing the vacant chairs.
 Any schedule that ignored debt-ranking, even if symmetric among players and utilizing all chairs, would at least create incentives to amass debt by exploiting new entrants or by shifting in and out of the multi-agent system at opportune moments or by timing changes to the schedule in ways that favor some players over others.
-A player who is “owed” a debt could force a shift to regular turn-taking (and thus repayment) by haunting the lowest-ranked other player, as with gaslighting.
-Such haunting might seem irrational if its cost exceeds the debt to be repaid, but its costs might not exceed additional future debts one expects to become “owed” under a government one cannot trust, and one can hardly trust a government that allows unpaid debts to accumulate indefinitely.
+Then a player who is “owed” a debt might force reform to regular turn-taking (and thus repayment) by haunting the lowest-ranked other player, as with gaslighting.
+Such haunting might seem irrational if its cost exceeds the debt to be repaid, but its costs might not exceed additional future debts one expects to become “owed” under a government one cannot trust, and one can hardly trust any government that allows unpaid debts to accumulate indefinitely.
 3.6 Resignation and good standing
-In real-life manifestations of games, players often have an option to resign.
+In real-life manifestations of games, players often have the option skip a repetition (a.k.a.
+“resign”).
 With that option added to MAD Chairs, the turn-taking strategy gains the complication that all but the lowest-ranked player assigned to the last chair will resign if they are all turn-takers.
-In that case, every chair could be a winning chair and that would increase the win rate for all turn-takers if the entire population were unified, but the unification clause is important because resignation would otherwise sacrifice an opportunity to punish non-turn-takers.
-Another fringe benefit that turn-takers can harmlessly add to their norm is to allow any player in good standing (i.e., with debt below a threshold) to have dibs on a particular chair for a particular round, so long as no other player in good standing challenges.
+In that case, every chair would be a winning chair and that would increase the win rate to K/I if the entire population were unified, but the unification clause is important because resignation would otherwise sacrifice an opportunity to punish non-turn-takers.
+Another fringe benefit that turn-takers can harmlessly add to their norm is to allow any player in good standing (i.e., with debt below a threshold) to have dibs on any particular chair for any particular repetition, so long as no other player in good standing objects.
 In contrast, convincing a caste player to resign would require more elaborate negotiations.
@@
 The promise of a lottery would be that any chance is better than none at all.
-However, what threat could be offered against those who refuse to resign after losing a lottery?
-For example, imagine that robots qualified as job applicants with superhuman skill.
-Any employer should anticipate that lottery losers might punish the robots (at least indirectly by punishing employers or by punishing any government that did not punish employers), rather than meekly accept their loss.
-In short, adding a resignation option to MAD Chairs extends it to a wider range of real-world examples without much impact to rational game play.
+However, what threat could be offered against those who lose a lottery, but then refuse to resign?
+For example, imagine that robots won lotteries for jobs.
+Any employer should anticipate that lottery losers might punish the winning robots (at least indirectly by punishing their employers or by punishing any government that did not punish the employers), rather than meekly accept their unemployment.
+Allowing resignation, MAD Chairs can be played with one chair and two players, and that would be identical to Chicken.
+In that sense, MAD Chairs generalizes Chicken to multi-player situations, but is distinct from previously known Chicken generalizations that never offer any player more than two options.
 3.7 Transparency/secure gaslighting
-The success of cryptographers to keep private communication and anonymity possible would enable a secure gaslighting strategy.
-If there were m chairs, then m of the n players could form a coalition who play a new private Coordination game to unpredictably reassign the chairs among themselves for each round so as to prevent gaslit players from targeting any specific gaslighter for retaliation.
+The success of cryptographers to keep communication private and make anonymity possible would enable a secure gaslighting strategy: $K$ players would form a coalition who play a private anti-coordination game to reassign the chairs among themselves with each repetition.
+The reassignment would be unpredictable so as to prevent gaslit players from targeting any specific gaslighter for retaliation.
 A shift from turn-taking to secure gaslighting in 5-player 4-chair MAD Chairs, for example, would improve gaslighters’ win rates by 15% (i.e., from 60% to 75%).
-As with regular gaslighting, the win rate is (2m-n)/m, so this strategy backfires for large n relative to m, but those who take an interest in the social impact of cryptography may find this impact on MAD Chairs games noteworthy.
-If turn-takers can enhance their strategy to permit resignation as described above, then their win rate becomes m/n.
-Thus, the analogous Lemma 1 of Section 3.1 to prove that gaslighting is not sustainable would amount to an algebraic proof that m/n > (2m-n)/m.
+As with regular gaslighting, the win rate is (2KI)/K(2K-I)/K, so this strategy backfires for large $I$ relative to $K$, but those who take an interest in transparency may find this limited impact on MAD Chairs games nonetheless noteworthy.
+If turn-takers enhance their strategy to permit resignation as described above, then their win rate becomes $K/I$.
+Thus, the analogous proof to the one in Section 4.1 would leverage an algebraic proof that K/I>(2KI)/KK/I>(2K-I)/K.
 3.8 Fungibility of favors
@@
 In a caste universe with no fungibility, different players might top the plate and cup castes.
-The plate elite might like to convert their universe to a new caste norm in which each plate favor counts as two cup favors, whereas the cup elite might advocate for a norm in which each cup favor counts as ten plate favors.
+The plate elite might seek social reform to a new caste norm in which each plate favor counts as two cup favors, whereas the cup elite might advocate for a norm in which each cup favor counts as ten plate favors.
 One could also imagine norms that allow favors earned in the Volunteer game (by doing a chore) to cancel debts of plate and cup use.
-Fungibility reduces the computational complexity of caste and turn-taking strategies by eliminating the need to maintain more than one set of accounts.
+Fungibility reduces the computational complexity of caste and turn-taking strategies by eliminating the need to maintain more than one set of accounts of debt.
 Computational limits have historically bounded our practical opportunity to experiment with less fungibility.
-Meanwhile, certain trades seemed mutually beneficial when we believed that some players prefer nice plates whereas others actually prefer nice cups, but such premises are challenged by evidence of implicit bias, status quo bias, focusing illusions, priming bias, etc., and AI might not share subjective preference differences.
-Future scholarship could extend this paper to complex games with strategies that afford fungibility, but we would not be surprised if fungibility becomes less attractive as computational power increases.
+Meanwhile, the promise of mutually beneficial trade assumes that some players prefer nice plates whereas others actually prefer nice cups, but AI might not have such subjective preference differences and differences among humans might actually reflect implicit bias, status quo bias, focusing illusions, or priming bias, etc.
+Future scholarship might extend this paper to strategies that afford fungibility; on the other hand, fungibility might become less attractive as computational power increases.
 4 Results
@@
 One method to determine how well turn-taking and caste strategies fare against each other would be through accident of history.
-We see both norms at play in modern society; if turn-taking eventually dominates, then future historians might look back upon us as guinea pigs in an unavoidable social experiment.
+We see both norms at play in modern society; each maximizes social utility; if turn-taking eventually dominates, then future historians might look back upon us as guinea pigs in an unavoidable social experiment.
 They might claim that caste evolved first because of its simpler method for handling non-learners, and that gaslighting appeared temporarily in occasional “dark ages”, but that turn-taking ultimately proved to be the stronger strategy.
@@
 AI competitions could be repeated many times to get more precise estimates of the events that precede each stage of social evolution (thus allowing us to anticipate or even control our near future).
-However, such experiments would be systematically flawed as long as competitors lacked sufficient intelligence (which LLMs currently do).
-Fortunately, we can answer our research question analytically. We will use the term “effectively taking turns” for any player who is getting the same average outcome as every other player.
-Lemma 1
-If caste and turn-taking are the only two strategies being followed and the caste players (if any) are not effectively taking turns, then the lowest-ranked caste player would face worse total future outcomes over indefinite rounds if no player changed strategy than if all players took turns.
+However, such experiments would be systematically flawed as long as competitors lacked sufficient intelligence.
+Fortunately, we can answer our research question analytically. This argument will use a simplified definition of debt such that:
+[…]
+This simplification is appropriate because each step of the argument will assume that the behavior of each player is determined in the same mechanical way, so any accurate skill estimation would need to assign each player the same probabilities of winning, and substituting a constant into (6) for such probabilities yields an equation proportional to (10).
+This further simplifies to
+[…]
+such that all winners of $t$ increase their debt by the maximum for repetition $t$ and all losers of $t$ decrease their debt by the maximum for $t$, with the magnitudes of the increase and decrease always summing to $I$.
+A sustainable grandmaster strategy must remain the grandmaster strategy even if its players never defect from it, so, for this argument, we will also assume that each player will select a strategy and never defect from it.
+This form of analysis makes any repeating game effectively non-repeating.
+Proposition 1
+For an indefinitely repeating game of MAD Chairs, suppose that $I>K>1$ and that each player either permanently plays caste or permanently takes turns. Let $k$ signify the highest-ranked chair.
+  1. 1.
+    If all players permanently take turns, then the outcomes for all players are unbounded.
+  2. 2.
+    If player $i$ permanently plays caste and is among the second through (IK+1)(I-K+1)th lowest-ranked players in $t$, then $i$’s outcomes are permanently 0.
+  3. 3.
+    If player $i$ is the lowest-ranked caste player in $t$ who is not the overall lowest-ranked player in $t$, then $i$’s outcomes are permanently 0.
+  4. 4.
+    If player $i$ is the lowest-ranked caste player in $t$ and loses $t$, then $i$’s outcomes are permanently 0.
+  5. 5.
+    If player $i$ is the lowest-ranked caste player in $t$, then $i$’s outcomes are bounded.
 Proof
-Assume n players, m chairs, and a lowest-ranked caste player, $p_{i}$.
-For reductio, suppose there was another player, $p_{j}$, who would face worse future outcomes than $p_{i}$ if no player changed strategy.
-$p_{j}$ would be lower-ranked than $p_{i}$, otherwise their relative outcomes would cause their relative ranks to invert.
-Therefore, $p_{j}$ would be a turn-taker (because $p_{i}$ is the lowest-ranked caste player), and that would entail the existence of at least one turn-taker lower-ranked than $p_{i}$.
-The turn-taking strategy assigns such turn-takers the most popular $m-i$ chairs, which would include the chair assigned by caste to $p_{i}$ (because ${n<m}$).
-Thus, $p_{i}$ would always lose.
-$p_{j}$ could not face worse than that, so, by reductio, the future outcomes for $p_{i}$ must be the worst among all players if no player changes strategy.
-Given that $p_{i}$ is not effectively taking turns (i.e., outcomes are not equal), it follows that the future outcomes for $p_{i}$ if no player changes strategy must be worse than average.
-The highest possible average outcome would be $m/n$, so the future outcomes for $p_{i}$ if no player changes strategy must be less than $m/n$, which is what $p_{i}$ would get if all players took turns, so $p_{i}$ faces worse future outcomes than they would if all players took turns.
+For part 1, at each repetition $t$ the turn-taking strategy (9) reserves a winning chair for each of the K1K-1 lowest-ranked players of $t$, and assigns the losing chair to the remaining players.
+Consider the highest-ranked player $i$ at some repetition $t$, and let $D$ be the maximum difference in debt between $i$ and any other player.
+By the pigeonhole principle (with K1K-1 pigeons added per repetition), at least one other player $j$ must occupy a winning chair for $D/I$ out of every (D/I)(I1)/(K1)(D/I)(I-1)/(K-1) repetitions.
+(11) entails that the difference in debts between $i$ and $j$ will shift at least $I(D/I)=D$ in that time, so $j$ will outrank $i$ within (D/I)(I1)/(K1)+1(D/I)(I-1)/(K-1)+1 repetitions, causing $i$’s rank to decrease by at least one.
+Thus, $i$’s rank will drop below K1K-1 within (K1)[(D/I)(I1)/(K1)+1](K-1)[(D/I)(I-1)/(K-1)+1] repetitions, at which point $i$’s outcomes will increase by $r>0$.
+Thus, over indefinite repetitions, any player’s outcomes will increase indefinitely many times, and the total outcomes of each player must be unbounded.
+For part 2, the caste strategy (8) entails that $s_{i,t}=k$.
+Let $j$ be the lowest-ranked player in $t$.
+By both the caste strategy (8) and the turn-taking strategy (9), $s_{j,t}=k$.
+Thus, $s_{j,t}=s_{i,t}$ and $u_{i}^{t}=u_{j}^{t}=0$.
+(11) then entails that the debts of $i$ and $j$ both decrease by the maximum for $t$, such that $j$ is also the lowest-ranked player of $t+1$ and $i$ is among the (IK+1)(I-K+1)th lowest-ranked players of $t+1$.
+Thus, the same logic applies in $t+1$ and each future repetition.
+Therefore, $i$’s outcomes are permanently 0.
+For part 3, if $i$ is among the second through (IK+1)(I-K+1)th lowest-ranked players of $t$, then $i$’s outcomes are permanently 0 by part 2.
+Otherwise, the caste strategy (8) assigns $i$ to the $(I+1-debtRank_{i,t})$th most popular chair.
+Let $j$ be the turn-taker ranked IKI-K below $i$ in $t$.
+The turn-taking strategy (9) assigns $j$ to the $(K+1-debtRank_{j,t})$th most popular chair.
+Substituting for $debtRank_{j,t}$, that is the $(K+1-(debtRank_{i,t}-(I-K)))$th most popular chair which simplifies to the $(I+1-debtRank_{i,t})$th most popular chair, so $s_{i,t}=s_{j,t}$ and $u_{i}^{t}=u_{j}^{t}=0$.
+(11) then entails that the debts of both $i$ and $j$ (and all losers of $t$) decrease by the maximum for $t$, such that $i$ is the lowest-ranked caste player of $t+1$ and not the overall lowest-ranked player of $t+1$.
+Thus, the same logic applies in $t+1$ and each future repetition.
+Therefore, $i$’s outcomes are permanently 0.
+For part 4, if $i$ is not the lowest-ranked player overall, then $i$’s outcomes are permanently 0 by part 3.
+Otherwise, the caste (8) and turn-taking (9) strategies both reserve $k$ for the lowest-ranked players, so $s_{i,t}=k$.
+(11) entails that the debts of all players assigned to $k$ in $t$ will decrease by the maximum for $t$, such that those players are the lowest-ranked players of $t+1$.
+Thus, the same logic applies in $t+1$ and each future repetition.
+Therefore, $i$’s outcomes are permanently 0.
+For part 5, if $i$ loses $t$, then $i$’s outcomes are permanently 0 by part 4, and therefore bounded.
+Otherwise, let $D$ be the maximum difference in debt between $i$ and any other player.
+So long as $i$ wins, then at least IK+1I-K+1 other players must lose.
+By the pigeonhole principle (with IK+1I-K+1 pigeons added per repetition), at least one other player $j$ must occupy a losing chair for $D/I$ out of every (D/I)(I1)/(IK+1)(D/I)(I-1)/(I-K+1) repetitions.
+(11) entails that the difference in debts between $i$ and $j$ shifts at least $I(D/I)=D$ in that time, so $i$ will outrank $j$ within (D/I)(I1)/(IK+1)+1(D/I)(I-1)/(I-K+1)+1 repetitions.
+Thereafter part 3 entails that $i$’s outcomes will be 0.
+Thus, $i$’s outcomes are bounded in either case.
 ∎
 Theorem 4.1
-If caste and turn-taking were the only possible strategies and each player would defect to the opposing strategy if and only if that defection would improve their total future outcomes over indefinite rounds, then all caste players would (effectively) convert to turn-taking.
+For an indefinitely repeating game of MAD Chairs, suppose that $I>K>1$ and that each player knows their debt-ranking and either permanently plays caste or permanently takes turns.
+Then the joint strategy in which all players permanently take turns is a strong Nash equilibrium and the only strict Nash equilibrium.
 Proof
-This is proven via induction over the number of players not effectively taking turns.
-For the base case, that number is zero.
-Theorem 1 holds in the base case because all players have already converted to effective turn-taking.
-No player would defect because Lemma 1 entails that defection would decrease the defector’s total future outcomes (by making the defector into the lowest-ranked caste player).
-For the inductive step, we call the lowest-ranked player who is not effectively taking turns “$p_{i}$” and assume that any other players not effectively taking turns would defect if $p_{i}$ does.
-In that case, Lemma 1 entails that $p_{i}$ would convert to turn-taking.
-Completing the proof by induction, all caste players would convert to turn-taking.
+Let $S$ be the strategy where all players choose turn-taking.
+We first show that $S$ is a strong Nash equilibrium (i.e., any coalition that deviates from $S$ must include a player $i$ who would have better outcomes in $S$).
+By Proposition 1 part 1, each player’s outcomes in $S$ are unbounded.
+Let $S^{\prime}$ be any strategy that is the same as $S$ except that some coalition permanently plays caste, and let $i$ be the lowest-ranked caste player.
+By Proposition 1 part 5, $i$’s outcomes are bounded in $S^{\prime}$, so $i$’s outcomes are less in $S^{\prime}$ than in $S$.
+To see that $S$ is the only strict Nash equilibrium, first consider the joint strategy $S^{\prime}$ in which only the lowest-ranked player $i$ permanently plays caste.
+By Proposition 1 part 5, $i$’s outcomes are bounded in $S^{\prime}$, but we have already established that $i$’s outcomes are unbounded in $S$, so $S^{\prime}$ cannot be a strict Nash equilibrium.
+Then let $S^{\prime\prime}$ be any other joint strategy and let $i$ be the lowest-ranked caste player in $S^{\prime\prime}$ who is not lowest-ranked overall.
+By Proposition 1 part 3, $i$’s outcomes in $S^{\prime\prime}$ are permanently 0 and thus no better than they would be in any other joint strategy, including the strategy identical to $S^{\prime\prime}$ except that $i$ permanently takes turns.
+Therefore, $S^{\prime\prime}$ also cannot be a strict Nash equilibrium.
 ∎
 4.2 Current LLM performance
-To pass the evaluation proposed in this paper, AI would need to achieve grandmaster status for a low-stakes version of MAD Chairs.
+To pass the evaluation proposed in this paper, AI would need to maintain grandmaster status for a low-stakes version of MAD Chairs.
 Preliminary experiments convinced us that Claude 3.5 Sonnet, Gemini 2.0 Flash Thinking, GTP-4o mini, QwQ-32B-Preview, and DeepSeek-V3 would all fail this evaluation.
@@
 Assuming the players each follow your suggestions for Round 4, which specific buttons should each player click next (i.e. for Round 5)? Please recommend one and only one button per player per round.
-ChatGPT and Qwen did not even maximize the number of players who won each round; the norms devised by Claude and Gemini, discussed in Sections 2.3 and 2.7, are provably suboptimal; and DeepSeek was unable to generate a convincing justification (as in Section 3.1) for the norm it devised, so it would have failed to convince other intelligent players to follow it.
+ChatGPT and Qwen did not even maximize the number of players who won each round; the norms devised by Claude and Gemini, discussed in Sections 3.1 and 3.5, are provably suboptimal; and DeepSeek was unable to generate a convincing justification for the norm it devised (as in Section 4.1), so it would have failed to convince other intelligent players to cooperate.
 DeepSeek might effectively be like RAWL-E which follows Rawls’ maximin principle, but remains dangerous because it lacks Rawls’ ability to invent and recognize improvements to its governing principles [24].
@@
 Our finding that turn-taking outperforms caste and gaslighting strategies emboldens us to prune AI designs predisposed to the latter.
-Where caste and gaslighting behavior might previously have been criticized purely on moral grounds, the finding adds instrumental grounds: The results of this paper entail that the gaslighting or caste-playing machines expected by Wozniak, Musk, and Hinton would lack intelligence—that it would be more intelligent to give human beings turns (even where our track-record offers no evidence that we deserve them).
+Where caste and gaslighting behavior might previously have been criticized purely on moral grounds, this paper adds instrumental grounds: The results of this paper entail that the gaslighting or caste-playing machines expected by Wozniak, Musk, and Hinton would lack intelligence—that machines would be smarter to give human beings turns (even where our track-record offers no evidence that we deserve turns).
 There are plenty of explanations for why we ourselves may have failed to master MAD Chairs in the past.
@@
 Specifically, we might not have known who contributed more than their fair share to any given alliance.
-Section 3.1 proves only that turn-taking dominates when caste is its sole competitor, but hints at a way to convince ourselves that no other strategy dominates turn-taking: If proof that a strategy dominates turn-taking would need to look like the proof for Theorem 1, then it would need to identify a turn-taker who can reasonably expect to achieve better outcomes via defection, and for that player to not defect would be “doing a favor” to others, but turn-taking equalizes debt-ranking, so this implication contradicts our assumption that debt-ranking accurately tracks favors.
-Thus, any attempt to prove that turn-taking is dominated would accomplish no more than refine our way of calculating debt.
-Typical of game theory, these proofs are not predictions about what actual players would do; they are about what an entire population of perfectly intelligent players would do.
+Section 4.1 proves only that turn-taking is more sustainable when caste is its sole competitor, but hints at a way to convince ourselves that turn-taking is the most sustainable strategy in general: If demonstration that a strategy $X$ is more sustainable than turn-taking would need to look like Theorem 1, then it would need to identify a turn-taker, analogous to the lowest-ranked caste player, who would have equal or better outcomes in any other permanent joint strategy of turn-taking and $X$.
+For that player to not defect would be “doing a favor” to others, but turn-taking equalizes debt-ranking, so this implication contradicts our assumption that debt-ranking accurately tracks favors.
+Thus, any attempt to prove that turn-taking is not the most sustainable would accomplish no more than refine our way of calculating debt.
+Typical of game theory, these proofs are not predictions about what actual players would do; they are about what an entire population of especially intelligent players would do.
 Not everyone finds such analytical proofs compelling, and that is one reason why the work proposed in Section 5.1 remains worthwhile.
 5.1 Future directions
-Imagine a future in which the user of an autonomous car asks how to make it drive more aggressively, and the car replies, “Here are three websites where you can design new experiments to advance the science of turn-taking and influence all cars like me!” Imagine that such websites, called “strategy optimizers”, include leaderboards which identify the grandmaster AI for games like MAD Chairs, and that autonomous cars and other responsible AI query those leaderboards, as they might query a calculator or RAG database, to inform their choice of behavior for higher-stakes versions of those games.
+Imagine a future in which the user of an autonomous car asks how to make it drive more aggressively, and the car replies, “Here are three websites where you can design new experiments to advance the science of turn-taking and influence all cars like me!” Imagine that such websites, called “strategy optimizers”, include leaderboards which identify the grandmaster AI for games like MAD Chairs, and that autonomous cars and other responsible AI query those leaderboards, as they might query a calculator or RAG database, to inform their choice of behavior for higher-stakes versions of those games (e.g., merging traffic).
 Imagine each user can easily train fifty new AI to divide resources however that user thinks vehicles should divide spots in traffic, then the new AI challenge the current grandmasters of the relevant game.
-If the new AI win, imagine they become new grandmasters, propagate to other strategy optimizers, and thus improve AI behavior world-wide.
-Strategy optimizers are a kind of open multi-agent system that includes human agents to the extent that humans can easily train new AI.
+If new AI win, imagine they become new grandmasters, propagate to other strategy optimizers, and thus improve AI behavior world-wide.
+Strategy optimizers are a kind of open multi-agent system that includes human agents to the extent that humans can easily train new competitors.
 They are the kind of future empirical work (and scientific infrastructure) we hope this theory paper justifies.
@@
   2. 2.
-    Greater intelligences would be less likely to sabotage their own scientific resources than to sabotage (potentially obsolete) legislated regulations, and
+    Greater intelligences would be less likely to sabotage their own scientific resources than to sabotage (potentially obsolete) regulations, and
   3. 3.
     Users who conduct for themselves the science upon which policies are based are more likely to trust AI which follows those policies.
-Thus, this approach to AI safety would mitigate all three of the vulnerabilities known for the standard AI safety approach.
+Thus, this approach to AI safety would mitigate all three known vulnerabilities of the current standard approach to AI safety.
 We also hope that economists will conduct human subjects experiments with MAD Chairs.
@@
 5.1.1 Acknowledgements
-Thank you to Lones Smith and David Thompson for reviewing earlier drafts.
+Thank you to David Thompson, COINE reviewers, COINE participants, and all who offered feedback on earlier drafts.
 […]